### v2

# Logical computation demonstrated with a neutral atom quantum processor

Abstract
Transitioning from quantum computation on physical qubits to quantum computation on encoded, logical qubits can improve the error rate of operations, and will be essential for realizing valuable quantum computational advantages. Using a neutral atom quantum processor with $256$ qubits, each an individual Ytterbium atom, we demonstrate the entanglement of $24$ logical qubits using the distance-two $\llbracket 4,2,2\rrbracket$ code, simultaneously detecting errors and correcting for lost qubits. We also implement the Bernstein-Vazirani algorithm with up to $28$ logical qubits encoded in the $\llbracket 4,1,2\rrbracket$ code, showing better-than-physical error rates. We demonstrate fault-tolerant quantum computation in our approach, guided by the proposal of Gottesman (2016), by performing repeated loss correction for both structured and random circuits encoded in the $\llbracket 4,2,2\rrbracket$ code. Finally, since distance-two codes can correct qubit loss, but not other errors, we show repeated loss and error correction using the distance-three $\llbracket 9,1,3\rrbracket$ Bacon-Shor code. These results begin to clear a path for achieving scientific quantum advantage with a programmable neutral atom quantum processor.
I Introduction
Quantum computers can enable faster, more accurate solutions than classical computers. However, the biggest challenge to scaling up quantum computers is errors, or “noise.” Physical qubits and their operation are inherently noisy. State-of-the-art two-qubit gate error rates depend on the qubit implementation, and in some flagship quantum processors range, e.g., from $p=0.1\%$ for an ion trap processor MBA+ (23) to an average of $p=0.4\%$ for a superconducting qubit processor GC (24). If a circuit cannot tolerate errors, it can only run for roughly $1/p$ steps, or fewer, which is not nearly enough to achieve large-scale, advantageous computation. Even the best physical qubits are insufficiently reliable, without error correction, to run even moderate-size quantum algorithms.
Indeed, obtaining a quantum advantage in scientific applications will likely require at least $100$ qubits with error rate $p$ at most $10^{-6}$, while a quantum advantage for industrial-scale problems may require upwards of $1000$ qubits with $p\sim 10^{-12}$ BMT+ (22). Such error rates will not be achievable at the physical level. Fortunately, encoding qubits in quantum error-correcting codes, thereby creating “logical” qubits, should allow dramatic reductions in effective error rates. The operations must also be encoded, according to fault-tolerant constructions AGP (06). Thus, by using quantum error correction and fault tolerance, noise in quantum hardware can be addressed through software, albeit with a cost in physical qubit overhead.
In this work, we present a step toward achieving scientific quantum advantage by demonstrating reliable quantum computation using logical qubits that outperforms corresponding physical-qubit based computation. We present our results on a neutral atom quantum processor. Neutral atom quantum architectures recently have been in rapid development. Previously, neutral atom architectures were predominantly analog in nature. Recent progress has enabled neutral atom approaches to transition to digital, gate-based computation EBK+ (23); SST+ (23); MLP+ (23); RCC+ (24); FTS+ (24); MSS+ (24). The promise of neutral atoms as qubits is their configurability, both in count and connectivity, while also maintaining low error rates. The neutral atom quantum processor we use in this work has up to 256 qubits (Ytterbium atoms), with all-to-all connectivity enabled by atom movement, along with high two-qubit physical gate fidelity and atom loss detection. This allows for flexible computation, as well as the ability to both use and optimize over a wide range of quantum error-correcting codes, including those requiring non-local connectivity, for the underlying neutral atom architecture. We refer to the system for detecting and correcting errors, also during computation, as the qubit virtualization system, as encoded, logical qubits are virtual qubits made up of some number of physical qubits.
Through co-optimization of fault-tolerant quantum error correction protocols within the qubit virtualization system and the neutral atom quantum processor design, we demonstrate the creation of logical qubits, their logical entanglement, and logical computation, while correcting for lost atoms.
Figure 1: Summary of results for the $24$-qubit cat state. Shown is the error for measurements in the $X$ basis, the error for measurements in the $Z$ basis, and their sum, for three cases. A total less than $50\%$ implies entanglement. The baseline (left) is an unencoded $24$-qubit cat state; the error bars are large because we have conservatively chosen only the best run instead of averaging all baseline runs, which would further increase the separation between logical and physical performance. See Fig. 7 for more details. The encoded bars are based on different ways of interpreting the same data for $\llbracket 4,2,2\rrbracket$-encoded cat states. In one case (middle), we reject trials with a detected error, but attempt to correct for lost qubits. In the other case (right), we reject trials with any lost qubit or detected error.
Figure 2: Summary of results for the Bernstein-Vazirani algorithm with an $n$-bit secret string. The string is always $1^{n}$, as the corresponding oracle is the most difficult to implement. The success probability is the probability of measuring $1^{n}$, conditioned on acceptance. For the unencoded algorithm (black), we accept when the $n$ measured atoms are present. For the encoded algorithm (red), we accept when all of the $n$ measured code blocks are decodable; in particular, it can detect faults and correct for lost atoms. The lighter points are based only on those trials with no initial atom loss.
Key results. We demonstrate a variety of reliable quantum computations that outperform physical qubit computation using logical qubits co-designed with a programmable neutral atom quantum processor. First, we show entanglement of $24$ logical qubits in a “cat” state (Fig. 1). This is the largest encoded cat state on record to date, illustrating the scale available in neutral atom processors. We then demonstrate quantum algorithms encoded into logical qubit computations. We implement the Bernstein-Vazirani algorithm BV (97); CEMM (98), a quantum algorithm well known for showing the higher efficiency of quantum computers versus classical. Our implementations involve as many as $28$ logical qubits, encoded in $112$ physical qubits (Fig. 2). Next, we demonstrate fault tolerance guided by a proposal of Gottesman Got (16). These experiments incorporate repeated error detection and, notably, repeated loss correction, a critical feature for scaling neutral atom processors. Each of the above experiments are encoded to distance two; they are capable of detecting phase and bit-flip errors and correcting atom loss. Finally, we demonstrate correcting both loss and errors, with repeated error correction in the distance-three Bacon-Shor code Bac (06).
Related work. Creating and verifying large cat states, also known as GHZ states GHZ (89), is a common benchmark for measuring hardware progress Kre (24). The quality of an experiment is typically expressed by the fidelity with the ideal cat state, and where a fidelity greater than $1/2$ is the success criterion. For example, Moses et al. MBA+ (23) prepared the $20$-qubit cat state with an $86\%$ fidelity, and the $32$-qubit cat state with an $82\%$ fidelity, on the Quantinuum H2 trapped-ion processor. Using superconducting qubits, Bao et al. BXS+ (24) prepared a $60$-qubit cat state with a $59\%$ fidelity. For encoded qubits, the published record to date is an encoded $12$-qubit cat state with at least a $99.8\%$ fidelity, also on H2 RAC+ (24). On a neutral atom system, Bluvstein et al. BEG+ (24) have prepared a four-qubit encoded cat state with a fidelity of $72(2)\%$ using error correction, ranging to $99.85^{+0.1}_{-1.0}\%$ using full error detection and therefore with much lower acceptance probability. Table 1 summarizes results with encoded cat states.
Table 1: Experiments preparing encoded cat states.
  Logical
Reference qubits Code Fidelity
HDHL (24) 4 $\llbracket 25,4,3\rrbracket$ $99.5^{+0.2}_{-0.4}\%$ to $99.7^{+0.2}_{-0.3}\%$
BEG+ (24) 4 $\llbracket 7,1,3\rrbracket$
[…]
72(2)error correction
$99.85^{+0.1}_{-1.0}\%$ error detection
RAC+ (24) 12 $\llbracket 16,4,4\rrbracket$ $99.82^{+0.12}_{-0.4}\%$ to $99.90^{+0.1}_{-0.3}\%$
This work 24 $\llbracket 4,2,2\rrbracket$
[…]
$73.4^{+3.1}_{-3.3}\%$ to $85.6^{+2.3}_{-2.5}\%$
error detection
and loss correction
$89.8^{+4.7}_{-6.6}\%$ to $91.5^{+4.3}_{-6.2}\%$
error and loss detection
The Bernstein-Vazirani algorithm for finding a secret string is another benchmark for quantum processors LGM+ (24). It has been implemented with unencoded qubits on superconducting processors RHK+ (20); DH (21); PL (23), ion trap processors FHM+ (16); DLF+ (16); WBD+ (19), and both processor types LMRM (17); MLM+ (19); BWM (21); LJV+ (23), for various lengths of the secret string, up to length $26$ for superconducting qubits PL (23) and up to length $20$ for ion traps LJV+ (23). It is difficult to compare these results fairly. Some test identification of the secret $s=1^{n}$, since it is hardest to implement; the corresponding oracle requires the most two-qubit gates. Others test against random strings $s$ or average over different Hamming weights, and this is much easier, particularly on ion traps, comparable to finding $s=1^{n/2}$. The success probabilities also vary considerably, e.g., from an exponentially small value for random $s$ with length $26$ PL (23), to $70\%$ for random $s$ with length $20$ LJV+ (23).
A critical capability for enabling more complex, reliable computation in a quantum processor is repeated detection and correction of errors. The latest generations of quantum hardware have enabled a variety of such demonstrations within the past few years. Repeated error detection with distance-two codes has been demonstrated on ion traps, including three rounds on eight logical qubits SBA (24), and 15 rounds on four logical qubits YDKR (24). Demonstrations based on Gottesman’s protocol Got (16) for fault tolerance with the $\llbracket 4,2,2\rrbracket$ code, but without including repeated error detection, include WWJ+ (18); HF (19); Vui (18); LGL+ (17); CCNH (21).
Repeated error correction, with higher-distance codes, has also been demonstrated. For single logical qubits, demonstrations include those on trapped-ion processors RBL+ (21); PBP+ (24) and superconducting processors KLR+ (22); Goo (23); GC (24); PNH+ (24) each with varying numbers of rounds, e.g., up to nearly $36$ rounds of error correction ($250$ syndrome extraction cycles) on the $\llbracket 49,1,7\rrbracket$ surface code, with a $0.143(3)\%$ logical error rate per cycle GC (24). Repeated error correction on multiple qubits has been implemented, including three rounds on two logical qubits PSR+ (24), and five rounds on eight logical qubits RAC+ (24), each while beating error rates of physical baselines.
II Physical machine description
Figure 3: Diagram of experimental architecture. 171Yb atoms, which each encode a physical qubit in their ground-state nuclear spin, can occupy tweezer traps within either a register array or an interaction zone (IZ). Up to 8 pairs of atoms interact at a time in the IZ, driven by two-qubit (2Q) gate lasers to implement controlled phase gates on the atoms in each pair. Arbitrary single-qubit (1Q) operations can be applied in parallel to the atoms within a row of the register; here one atom is missing. Each atom has a home register location, indicated by dashed lines, where it is originally loaded and where 1Q gates can be applied. Movement of atoms between the two zones enables arbitrary connectivity for 2Q gates. Up to eight atoms from the same row can move in parallel, but their paths cannot cross.
All results presented in this work were obtained on a quantum processor based on reconfigurable arrays of neutral 171Yb atoms, depicted in Fig. 3, with the qubits encoded in the ground-nuclear spin states (1S0, $m_{F}=\pm 1/2$). Aspects of the system have previously been described in BBB+ (22); NC+ (23); NKC+ (24), as well as an accompanying publication MSS+ (24). Further system and benchmarking details will follow in a separate manuscript, in preparation.
By moving atoms between arrays of static optical tweezer traps using mobile tweezers, this platform enables arbitrary connectivity between qubits, for a large numbers of qubits. Single-qubit operations and qubit storage are performed in a register array, which is loaded in an iterative manner from a repeatedly filled reservoir NKC+ (24). Two-qubit gates are performed in a separate interaction zone (IZ). For this work, up to 256 atoms are available for computation, and up to eight atoms within the same row may be moved between the register and IZ at a time, subject to constraints associated with intermodulation in the acousto-optic deflectors used to generate the moving tweezers. Typical atom loss probabilities per move are below 0.1%. State-resolved, low-loss readout is performed by transferring atoms from the tweezer arrays into traps formed by cavity-enhanced optical lattices, which allows for state and atom detection with error and loss rates at the few parts-per-thousand level (see NKC+ (24) for a detailed characterization of state preparation and imaging). Importantly, the ability to perform state-selective, low-loss imaging allows us to distinguish between different types of error. An image performed prior to circuit execution allows the identification of sites successfully occupied by atoms, while a pair of images directly after circuit execution and after optical pumping within the qubit subspace allow determination of both the presence and state of the atoms.
In the encoded Bernstein-Vazirani experiment with $n=7$, for example, on average $0.52(7)\%$ of the $32$ atoms were not loaded, and $1054$ of $1200$ shots, or $85(2)\%$, had no initial loss. The fraction of qubits lost during the circuit averaged $2.3^{+0.2}_{-0.1}\%$. Initial loss varies, and more complicated circuits would have a higher circuit loss rate.
Single-qubit operations are performed in parallel on rows of atoms within the register, with the pulse area, phase, and presence of the gates controlled individually on each site using an architecture similar to that described in Ref. BBB+ (22). These gates are performed using two-photon Raman transitions driven by two independently controlled and site-resolved laser beams detuned roughly 10 GHz from the 3P1, $F=1/2$ manifold. We calibrate an $S_{X}$ gate on each site of the register, and perform virtual $Z$ rotations of arbitrary angle by changing the relative phase of the two drive lasers on subsequent gates. We characterize the performance of the single-qubit operations using a Clifford randomized benchmark sequence. At depths of up to 500 Clifford gates, we find an average fidelity of 99.85(2)per Clifford gate, averaged across three rows of the register spanning those used in this work. Because the nuclear spin qubit is well-isolated from differential light-shifts and has low sensitivity to magnetic fields, we do not automatically apply dynamical decoupling sequences for the circuits presented in this work.
Two-qubit controlled-phase (CZ) gates are performed in the IZ on up to eight pairs of atoms in parallel. These gates are mediated by excitation to a highly-excited Rydberg state ($65^{3}\mathrm{S}_{1}$). The sites of the IZ are spaced such that atoms occupying the same pair of sites are in the so-called “Rydberg blockade” limit, where the strength of interactions between the atoms exceeds other relevant energy scales, while atoms in different pairs interact weakly enough to be treated independently. All operations in the IZ are applied in parallel to all atoms within the IZ.
Atoms are excited to the Rydberg state using a sequential two-photon transition. State-selectivity is provided by the first stage of the two-photon transition, from 1S0, $m_{F}=-1/2$ to the metastable 3P0, $m_{F}=1/2$ state. Polarization selectivity, along with the narrow linewidth of the transition and slow Rabi rate (10 kHz) relative to the qubit frequency (376 kHz) ensure that only one qubit state is excited to the metastable state. From the metastable state, atoms are excited to the Rydberg state using a UV laser pulse (total duration 110 ns) that implements a gate similar to the time-optimal controlled-phase (CZ) gate described in JP (22). While the IZ traps are equally confining for the ground and metastable states, the register traps are near a tune-out wavelength for the metastable states, providing near-zero trapping potential. Thus, any gate errors that lead to population being left in the metastable state (such as some coherent population errors on either stage of the two-photon transition) are converted to loss and can be detected. This “erasure conversion” proves useful in identifying the location of errors WKPT (22).
We quantify the performance of our two-qubit gates using several variations on randomized benchmarking. In all presented cases, data is taken in parallel over the eight pairs of sites in the IZ, and averaged. To isolate the performance of the two-qubit gates themselves, in the absence of any errors that may be incurred through atomic movement, we perform interleaved randomized benchmarking (IRB) on static atoms within the IZ. In this protocol, we compare the return probability after a given number of random two-qubit Clifford operations (plus a return pulse calculated to return to a desired measurement state) to the same circuit with an additional CZ gate interleaved after each Clifford gate (and with an appropriately modified return pulse). By comparing the probability of returning to the measurement state versus the depth of the Clifford circuit with and without the added CZ gates, we isolate the error associated with the CZ gate alone, averaged over the two-qubit Hilbert space. Here, the atoms are kept in the IZ for all operations (unlike typical circuits, single qubit gates are applied in the IZ for this protocol). Fits to IRB data taken out to depth 34 indicate an infidelity per CZ gate of 0.56(15)and additional loss and population leakage outside of the qubit subspace of 0.27(6)%.
Table 2: Randomized benchmark infidelities, post-selected on an atom (1Q) or pair of atoms (2Q) remaining in the qubit subspace, and probability of one or both atoms being lost from the qubit subspace. Benchmarking sequences are described in the main text.
  Benchmark protocol Infidelity Loss + Leakage
1Q Clifford RB 0.15(2)0.007(1)%
2Q Static IRB 0.56(15)0.27(6)%
2Q Echoed RB with moves 0.4(1)0.5(1)%
2Q Echoed RB without moves 0.35(5)0.24(6)%
In actual circuits, atoms are moved between the register and IZ between single-qubit operations and two-qubit operations, to allow arbitrary connectivity. Because atom movement can cause heating, which may impact the performance of all following gates, IRB does not cleanly isolate the performance of the CZ gate when atoms are moved between arrays. For this, we rely on a variant of randomized benchmarking described in MSS+ (24) and similar to that in EBK+ (23), where concatenated circuit blocks each consist of a random single-qubit operation and two CZ gates separated by an echo pulse. A final gate sequence returns atoms to the measurement basis. We isolate the contribution from the two-qubit gates (along with any loss or infidelity due to associated movements, up to single-qubit phases which are canceled by the echo pulses) by subtracting the per-depth infidelity associated with the single-qubit operations. From this, we extract a per-CZ gate error of 0.4(1)and loss of 0.5(1)%. For comparison, the same procedure applied to atoms that remain static in the IZ for all operations yields a per-gate error rate of 0.35(5)and loss of 0.24(6)%. Table 2 summarizes the results of these benchmarks. The performance of our gates is sensitive to relative alignment of gate lasers and optical potentials, and other factors that may drift on the timescale of hours to days. The values presented here represent typical values over the time period in which the circuits presented in this work were run.
In our experiments, atom movement is scheduled using a combination of manual and algorithmic techniques, including the use of graph and annealing algorithms to minimize the number of atom movement operations.
III Four-qubit codes
The $\llbracket 4,2,2\rrbracket$ code VGW (96); GBP (97) is a stabilizer code Got (98) with stabilizer generators
$\begin{array}[]{cccc}X&X&X&X\\ Z&Z&Z&Z\end{array}$
and logical operators
$\begin{array}[]{rccccc}\overline{X}_{1}=&X&X&I&I&\hskip 22.40352pt{\color[rgb]{1,1,1}.}\\[-2.84544pt] \overline{Z}_{1}=&I&Z&I&Z\\ \overline{X}_{2}=&X&I&X&I\\[-2.84544pt] \overline{Z}_{2}=&I&I&Z&Z\end{array}$
The code can detect any single-qubit error; for example, a single-qubit $Y$ error flips both stabilizer generators. The code can correct for a single-qubit loss. For example, if the second qubit is lost and this is known (an erasure error), then logical qubit $1$ can still be decoded using the logical operators $\overline{X}_{1}\sim IIXX$, $\overline{Z}_{1}\sim ZIZI$, because, having multiplied by a stabilizer, these operators are not supported on the second qubit. However, once a qubit is lost, the code does not protect against errors on the remaining qubits, e.g., with the second qubit lost an $X_{1}$ qubit error results in a logical $\overline{X}_{1}$ error.
Observe that encoded ${|00\rangle}$ is the four-qubit cat state $\tfrac{1}{\sqrt{2}}({|0000\rangle}+{|1111\rangle})$. As shown in Figs. 4 and 5, an extra ancilla qubit is needed to prepare this state fault tolerantly Sho (96).
The $\llbracket 4,1,2\rrbracket$ code is actually the same code, but only using one of the two logical qubits. The other logical qubit is treated as a gauge qubit, meaning we can set it arbitrarily Bac (06). Unlike the $\llbracket 4,2,2\rrbracket$ code, the $\llbracket 4,1,2\rrbracket$ code can sometimes correct two qubit losses. For example, $\overline{X}_{1}$ can still be decoded if both qubits $1$ and $2$ are lost, or both $3$ and $4$, but not any other pair.
Between two $\llbracket 4,2,2\rrbracket$ or $\llbracket 4,1,2\rrbracket$ code blocks, transversal physical CNOT gates fault-tolerantly implement logical transversal CNOT. Other logical operations can also be executed fault tolerantly on a $\llbracket 4,2,2\rrbracket$ code block, including SWAP, CNOT, CZ and Hadamard ($H\otimes H$) on both qubits. Swapping the middle two physical qubits swaps the logical qubits. Swapping the first and third physical qubits implements a logical CNOT gate from $1$ to $2$. Applying the Hadamard gate $H$ transversally implements logical $H\otimes H$ and a logical swap. Letting $S_{Z}=\left(\begin{smallmatrix}1&0\\ 0&i\end{smallmatrix}\right)=\sqrt{Z}$, and $S_{X}=\tfrac{1}{\sqrt{2}}\left(\begin{smallmatrix}1&-i\\ -i&1\end{smallmatrix}\right)\propto\sqrt{X}$, observe that
$S_{Z}XS_{Z}^{\dagger}$ $=Y$ $S_{Z}YS_{Z}^{\dagger}$ $=-X$ $S_{Z}ZS_{Z}^{\dagger}$ $=Z$
$S_{X}XS_{X}^{\dagger}$ $=X$ $S_{X}YS_{X}^{\dagger}$ $=Z$ $S_{X}ZS_{X}^{\dagger}$ $=-Y.$
Simple stabilizer algebra then shows that applying $S_{Z}\otimes S_{Z}^{\dagger}\otimes S_{Z}^{\dagger}\otimes S_{Z}$ implements a logical CZ gate. Finally, using $H=S_{Z}S_{X}S_{Z}$, applying $S_{X}\otimes S_{X}^{\dagger}\otimes S_{X}^{\dagger}\otimes S_{X}$ implements a logical dual CZ gate, $(H\otimes H)\,\text{CZ}\,(H\otimes H)$.
IV Encoded $24$-qubit cat state
An $n$-qubit cat state can be written as $\tfrac{1}{\sqrt{2}}({|0^{n}\rangle}+{|1^{n}\rangle})$. Here, we prepare an encoded cat state on $24$ logical qubits, encoded with the $\llbracket 4,2,2\rrbracket$ code into $48$ physical qubits. The circuit we implement is explained in Fig. 4. Starting with a logical circuit to prepare $\tfrac{1}{\sqrt{2}}({|0^{24}\rangle}+{|1^{24}\rangle})$, we fault-tolerantly encode it into the $\llbracket 4,2,2\rrbracket$ code. We then arrange the qubits, four code blocks in each row with the ancilla qubits at the end, and schedule the gates and qubit movements as in Fig. 5. The unencoded baseline experiment prepares a $24$-qubit cat state using a separately optimized circuit, in Fig. 6.
[…]
Figure 5: A portion of the encoded circuit from Fig. 4 prepares an encoded four-qubit cat state. The gates in this circuit need to be scheduled. Two-qubit CZ gates are applied in the interaction zone, up to eight gates in parallel, and all the one-qubit gates on a particular qubit are applied at its initial location in the storage zone. One-qubit gates on a subset of qubits in the same row can be applied in parallel. Two atoms can be moved in parallel, but their paths cannot cross. Note that in this example, for clarity the two code blocks are arranged adjacent to each other in the storage zone, whereas in Fig. 4 they are interleaved.
Figure 6: Circuit to prepare a $24$-qubit cat state, unencoded. One-qubit gates are applied at an atom’s home register location.
The cat state has genuine multipartite entanglement, meaning that it is entangled across every bipartition of the $n$ qubits. Any state that has fidelity greater than $1/2$ with the ideal cat state also has genuine multipartite entanglement GS (10). However, states with $\leq 1/2$ fidelity may not even be entangled; the classical mixture $\tfrac{1}{2}({{|0^{n}\rangle}\!{\langle 0^{n}|}}+{{|1^{n}\rangle}\!{\langle 1^{n}|}})$ has fidelity $1/2$ with the ideal cat state, but no entanglement. Our experiments, as well as HDHL (24); RAC+ (24), measure $X$ basis and $Z$ basis error rates, $p_{X}$ and $p_{Z}$. The fidelity to the ideal cat state is between $1-p_{X}-p_{Z}$ and $1-\max(p_{X},p_{Z})$, essentially because possible $Y$ errors contribute to both $p_{X}$ and $p_{Z}$; this is the reason for the fidelity ranges in Table 1.
We find that for the encoded cat state, using error detection and loss correction, $p_{X}+p_{Z}=26.6^{+3.3}_{-3.1}\%$, while for the unencoded baseline $p_{X}+p_{Z}=42.0^{+11.3}_{-10.1}\%$. The results are summarized in Fig. 1 and Table 3. While the encoding is clearly beneficial, interestingly, it helps reduce $p_{X}$ much more than $p_{Z}$. The baseline error bar is large because rather than averaging over all trials, we chose for a challenging baseline the best of $27$ separate runs, each consisting of 300300300300400400400400 unencoded trials with $X$ and $Z$ basis measurement; Fig. 7 gives more details.
Figure 8 shows how conditioning on at most $k$ lost qubits, for various values of $k$, allows a smoother interpolation, in both acceptance probability and total error probability, between the different results in Table 3. Finally, Fig. 9 shows the different machine runs that combine to give the data in Table 3, plotted conditioned on no initial atom loss. The spread and correlations in the figure suggest some machine parameter drift.
Figure 7: Each trial of the unencoded baseline $24$-qubit cat state experiment included two cat states, one measured in the $X$ basis and the other measured in the $Z$ basis. Shown here are the baseline runs, each consisting of 300300300300400400400400 trials. Only cat states with no qubit loss are included, on average $135.3$ and $161.0$ trials for $X$ and $Z$ basis measurements, respectively. The high variance in the observed error rates complicates reporting a single baseline $X$ plus $Z$ basis total error rate. To be conservative, disadvantaging the encoded cases, we decided to report the best single run as a baseline. The red data point has the least total error. In the $X$ basis, $26$ out of $100$ cat states had a logical error, and in the $Z$ basis, $19$ out of $121$ cat states had a logical error. The baseline error bars in Fig. 1 are large because they are based on only this run, rather than on the average of all the runs in this plot.
Figure 8: Data from the encoded $24$-qubit cat state experiment, selecting only trials within specified loss criteria. In general, attempting to correct for more loss leads to a higher acceptance probability, but also a higher error rate. In each trial, an atom may be missing before the circuit even begins, a loading loss, or it may be lost during the circuit. The red points show the total $X$ basis plus $Z$ basis error rate, plotted against the average of the $X$ and $Z$ acceptance probabilities, based on either all 29 7602976029\,76029 760 trials, only those 17 0081700817\,00817 008 trials with no initial loss, or only those $858$ trials with no loss at all. For the blue points, we select only trials in which the total number of lost qubits, including load loss and circuit loss, is at most a certain threshold. Note that Fig. 9 only includes trials with no initial loss, and the acceptance probabilities in Fig. 9 are conditioned on no initial loss, so the normalization is different. (Figure 7 for the baseline only includes trials with no loss at all.)
Table 3: Data for the $24$-qubit cat state experiments. An experiment is accepted if it is neither prerejected nor postrejected. For the encoded experiments, a trial is “prerejected” if it is rejected during the initial state preparation because one of the ${|\overline{00}\rangle}=\tfrac{1}{\sqrt{2}}({|0^{4}\rangle}+{|1^{4}\rangle})$ states fails verification. A surviving trial is “postrejected” if a code block cannot be decoded, either because of a detected error or because of two or more lost qubits in the block. We have also filtered out the encoded data with no initial losses, and the data with no losses at all. For the unencoded baseline, a trial is prerejected if there are any initial losses and a surviving trial is postrejected if there are any circuit losses. As explained in Fig. 7, we have selected the best of the unencoded runs, in order to create a challenging baseline.
  Experiment Basis Trials Prerejected Postrejected $\Pr[\text{accept}]$ Errors Error rate
Cat-$24$ encoded: all data $X$ 14 4001440014\,40014 400 10 1591015910\,15910 159 $3026$ $8.4^{+0.5}_{-0.4}\%$ 148 $12(2)\%$
$Z$ 15 3601536015\,36015 360 11 0901109011\,09011 090 3283 $6.4(4)\%$ 170 $17(2)\%$
no initial loss $X$ $8433$ 5371 2122 $11.1(7)\%$ 115 $12(2)\%$
$Z$ $8575$ 5589 2185 $9.3(6)\%$ 115 $14^{+3}_{-2}\%$
no loss $X$ 395 66 221 $27^{+5}_{-4}\%$ 9 $8^{+6}_{-4}\%$
$Z$ 463 98 220 $31(4)\%$ 2 $1^{+3}_{-1}\%$
Cat-$24$ unencoded baseline $X$ 300 87 113 33(5)26 $26^{+8}_{-9}\%$
$Z$ 300 82 97 $40^{+6}_{-5}\%$ 19 $16^{+7}_{-6}\%$
(a) $X$ basis runs
(b) $Z$ basis runs
Figure 9: Encoded $24$-qubit cat state data from 45 $X$ basis measurement runs, and 48 $Z$ basis measurement runs. This data was taken overnight, while the machine ran unattended. Each blue point gives the acceptance probability, conditioned on no initial loss, and logical error rate for a run of, on average, $187$ trials for the $X$ basis data, and $179$ trials for the $Z$ basis data. (We attempted $320$ trials, but some trials failed to load all atoms.) For clarity, only some error bars are shown. The black squares are the averages over all the data, the numbers we report. We interpret the negative correlation between acceptance probability and logical error rate as evidence of some machine parameter drift between runs, despite automatic recalibrations.
V Encoded Bernstein-Vazirani algorithm
Consider a black box oracle $\{0,1\}^{n}\rightarrow\{0,1\}$, with a secret string $s\in\{0,1\}^{n}$, that on query $x$ returns $s\cdot x=s_{1}x_{1}\oplus\cdots\oplus s_{n}x_{n}$, with $\oplus$ denoting addition modulo $2$. Classically, $n$ queries to the oracle are required and sufficient to learn $s$. The Bernstein-Vazirani algorithm BV (97); CEMM (98) makes a single query to the quantum version of this oracle, $U_{s}=\sum_{x\in\{0,1\}^{n},b\in\{0,1\}}{{|b\oplus(s\cdot x)\rangle}\!{\langle b|}}\otimes{{|x\rangle}\!{\langle x|}}$, to learn $s$ through the following steps:
${|-,+^{n}\rangle}$ $\overset{U_{s}}{\longmapsto}\tfrac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}{(-1)^{s\cdot x}{|-,x\rangle}}$ $\overset{I\otimes H^{\otimes n}}{\longmapsto}{|-,s\rangle}\enspace.$
This classic algorithm was designed to prove an oracle separation between the classical randomized and quantum polynomial-time complexity classes $\mathsf{BPP}$ and $\mathsf{BQP}$.
[…]
Figure 10: The $n=19$ Bernstein-Vazirani algorithm involves $n+1=20$ logical qubits. In our experiment, we test the most challenging case, in which the secret string is $1^{n}$, so there are $n$ logical CNOT gates. Encoded into the $\llbracket 4,1,2\rrbracket$ code, each logical qubit becomes a block of four atoms. There are $4(n+1)$ atoms and $26n+$ 2 physical CZ gates. Four code blocks are arranged in each of the storage zone’s top five rows. The gates are scheduled so that the first block, once initialized, stays in the interaction zone. As transversal CZ gates are applied into block $j$, block $j+1$ is initialized.
The unencoded algorithm accepts if no measured atom is lost, and is successful if the measured string is $1^{n}$. The encoded algorithm tolerates losses, but rejects if any of the last $n$ code blocks is not decodable.
Figure 2 summarizes the results, with more details in Table 4. The encoded algorithms have a statistically significantly higher success probability than the unencoded baseline, albeit with a lower acceptance rate.
Table 4: Data for the Bernstein-Vazirani experiments. The encoded experiment has no preselection. A surviving trial is rejected if one of the last $n$ code blocks cannot be decoded, either because of a detected error or from too many lost qubits. (The first code block is ignored.) For the unencoded baselines, a trial is rejected if a measured qubit is lost.
      All data
$n$ Version Trials Rejected $\Pr[\text{accept}]$ Errors Error rate
7 encoded 1200 409 $66(3)\%$ 18 $2.3^{+1.2}_{-0.9}\%$
baseline 600 224 $63(4)\%$ 57 $15^{+4}_{-3}\%$
11 encoded 1200 547 $54(3)\%$ 22 $3^{+2}_{-1}\%$
baseline 600 196 $67(4)\%$ 65 $16^{+4}_{-3}\%$
15 encoded 1200 740 $38(3)\%$ 40 $9^{+3}_{-2}\%$
baseline 600 243 $59(4)\%$ 73 $20(4)\%$
19 encoded 1800 1359 $25(2)\%$ 43 $11(3)\%$
baseline 800 457 $43(3)\%$ 90 $26^{+5}_{-4}\%$
23 encoded 2400 1993 $17^{+2}_{-1}\%$ 67 $16^{+4}_{-3}\%$
baseline 3160 2379 $25^{+2}_{-1}\%$ 212 $27(3)\%$
27 encoded 26 0402604026\,04026 040 23 7892378923\,78923 789 $8.6(3)\%$ 498 $(2)\%2$
baseline 15 6001560015\,60015 600 12 8771287712\,87712 877 $17.5(6)\%$ 857 $31(2)\%$
Trials with no initial loss
$n$ Version Trials Rejected $\Pr[\text{accept}]$ Errors Error rate
7 encoded 732 233 $68(3)\%$ 7 $1.4_{-0.8}^{+1.3}\%$
baseline 425 76 $82_{-4}^{+3}\%$ 30 $9(3)\%$
11 encoded 471 181 $62(4)\%$ 4 $1.4_{-1.0}^{+1.8}\%$
baseline 540 142 $74(4)\%$ 59 $15_{-3}^{+4}\%$
15 encoded 543 308 $43(4)\%$ 15 $6_{-3}^{+4}\%$
baseline 526 176 $67(4)\%$ 66 $19(4)\%$
19 encoded 537 369 $31(4)\%$ 16 $10_{-4}^{+5}\%$
baseline 623 289 $54(4)\%$ 81 $24_{-4}^{+5}\%$
23 encoded 800 647 $19(3)\%$ 25 $16_{-5}^{+6}\%$
baseline 2173 1409 $35(2)\%$ 195 $26(3)\%$
27 encoded 4981 4443 $10.8_{-0.8}^{+0.9}\%$ 104 $19(3)\%$
baseline 10 3581035810\,35810 358 7695 $25.7(8)\%$ 797 $30(2)\%$
Computing the success probability conditioned on acceptance makes sense if we treat the Bernstein-Vazirani algorithm as a generic quantum algorithm, as done above. However, it was originally proposed as a quantum query algorithm, in which minimizing the number of oracle queries was of key importance. In this light, we can look at the data and ask whether with only one oracle query allowed (one trial), the encoded circuit still beats the unencoded circuit. In short, the answer is probably not. In this case, neither the encoded nor unencoded algorithms can condition on acceptance, and with only one trial it requires guessing the secret string. In particular, with a lost unencoded qubit or an undecodable code block the probability of guessing the corresponding bit is only $1/2$, and the expected Hamming distance of the guess away from the secret string increases by $1/2$. As shown in Table 5, if you are only allowed to make one query to either the unencoded or encoded oracles, and if your goal is to maximize the probability of guessing the secret string $s$, then for $n\geq 11$ it is better to use the unencoded oracle. On the other hand, if your goal is to minimize the expected Hamming distance of your guess from $s$, then it is better in all cases to use the encoded oracle.
These statements are not contradictory. They hold because the failure modes of the encoded and unencoded algorithms are quite different, as shown in Fig. 11. Intuitively, an error on the ancilla qubit in the unencoded algorithm typically results in a string like $1^{n/2}0^{n/2}$ and contributes on average $n/2$ to the expected Hamming distance from $s$. On the other hand, an error on the ancilla code block in the encoded algorithm often results in a string like $1^{n/2}\star^{n/2}$, where $\star$ denotes an undecodeable code block where one must guess a bit of $s$; this contributes only an average of $n/4$ to the expected Hamming distance. Typical errors in the encoded algorithm are heralded, so they contribute less to the expected Hamming distance error metric.
One query to either quantum oracle is much better than one query to the classical oracle. One classical query can only return one bit of $s$, so the probability of guessing $s$ is $\Pr[\text{guess}]=1/2^{n-1}$ and the expected Hamming distance to $s$ is $\operatorname{\operatorname{E}}[\text{Hamming}]=\tfrac{n-1}{2}$.
Table 5: Performance of unencoded and encoded Bernstein-Vazirani algorithms with a single oracle query. $\Pr[\text{guess}]$ is the probability of guessing the secret string based on the measured outputs, and $\operatorname{\operatorname{E}}[\text{Hamming}]$ the expected distance of the guess away from the secret string. This analysis is based on all trials, including ones with initial loss. For a classical oracle, $\Pr[\text{guess}]=1/2^{n-1}$ and $\operatorname{\operatorname{E}}[\text{Hamming}]=\tfrac{n-1}{2}$.
  Unencoded baseline Encoded algorithm
$n$ $\Pr[\text{guess}]$ $\operatorname{\operatorname{E}}[\text{Hamming}]$ $\Pr[\text{guess}]$ $\operatorname{\operatorname{E}}[\text{Hamming}]$
7 67(4)0.94 75(2)0.48
11 69(4)1.04 66(3)0.90
15 61(4)1.58 51(3)1.45
19 49(3)2.44 38(2)2.29
23 35(2)3.42 30(2)2.97
27 28.2(7)4.12 18.5(5)4.05
[…]
VI Repeated loss correction and transversal operations for the $\llbracket 4,2,2\rrbracket$ code
VI.1 Repeated loss correction with CZ and dual CZ logical gates
The circuits considered so far have only used loss correction and error detection at the very end, when decoding the final measurements. However, it’s important to be able to detect and correct errors between computational steps of an algorithm, to enable deeper computation. To that end, we explore using loss correction in the middle of a larger circuit, demonstrating repeated fault-tolerant loss correction. In order to make the circuits more interesting, between rounds of loss correction we apply some of the fault-tolerant logical operations given in Sec. III.
Figure 12 gives a circuit for fault-tolerant error detection for the $\llbracket 4,2,2\rrbracket$ code. It measures the stabilizer generators $X^{\otimes 4}$ and $Z^{\otimes 4}$ in parallel, with one measurement outcome flagging the other for possible correlated errors. Furthermore, the circuit in Fig. 12(b) replaces two of the data qubits with fresh qubits, so the full data block can be refreshed every two rounds of error detection. This allows us to use the $\llbracket 4,2,2\rrbracket$ code to correct a lost qubit during a computation, not just at the very end as in our logical cat state procedure.
Figure 12: The top circuit (a) measures the parity of four qubits, i.e., the syndrome of $ZZZZ$. However, it is not fault tolerant; a single $Z$ fault on the ancilla can spread undetected to a weight-two $Z$ error. The next circuit (b) remedies this, and simultaneously measures $XXXX$, using a trick from Rei (20). Now the bad $Z$ fault will be detected, or “flagged,” by the $XXXX$ measurement, and similarly the $ZZZZ$ measurement will flag a bad $X$ fault. This circuit gives an efficient, fault-tolerant error detection procedure for the $\llbracket 4,2,2\rrbracket$ code. However, it does not correct for a lost qubit. The bottom circuit (c) fixes this, halfway, by switching the direction of the final two CNOT gates. (This is equivalent to inserting swap-based “leakage-detection units” CBO+ (24).) The last two data qubits are replaced with fresh ancillas, so in particular if one of them is lost it will be replaced. Thus we can replace all four qubits in two rounds of error detection. Future work will investigate this paradigm by combining mid-circuit measurement, qubit reuse and reinitialization, along with periodic atom cooling to remove a fundamental obstacle to neutral atom quantum computing by limiting the depth of circuit any atom must undergo.
We design and execute experiments that repeatedly apply logical CZ gates on a $\llbracket 4,2,2\rrbracket$ code block. Recall that $S_{Z}\otimes S_{Z}^{\dagger}\otimes S_{Z}^{\dagger}\otimes S_{Z}$ implements a logical CZ gate. Error detection with qubit refresh is inserted between each pair of CZ gates, as illustrated in Fig. 13. Note that with $j+1$ logical CZ gates, there are $j$ rounds of error detection, with $5+8j$ CZ gates applied to $5+2j$ qubits.
Figure 13: (a) Encoded circuit that applies two rounds of encoded CZ gates to ${|\overline{++}\rangle}$, encoded in the $\llbracket 4,2,2\rrbracket$ code, with two rounds of error detection. Together, the two error detection rounds refresh all four data qubits. (b) The corresponding logical circuit. Since the CZ gates cancel out, both measurements should give $+$. (c) The baseline circuit that we run for comparison. The $X$ echo pulses cancel out coherent dephasing errors from the CZ gates and movement. Both measurements should give --. This baseline performs better than applying the two CZ gates sequentially without leaving the interaction zone.
We also ran a very similar experiment, except applying logical dual CZ gates, $(H\otimes H)\,\text{CZ}\,(H\otimes H)$, using $S_{X}\otimes S_{X}^{\dagger}\otimes S_{X}^{\dagger}\otimes S_{X}$.
The encoded experiments can be compared against corresponding unencoded baselines. Given the system architecture, there are a few reasonable baselines from which to choose. For example, we may choose to return qubits to the storage zone between each CZ, or not. We have chosen to select unencoded baselines in which, as shown in Fig. 13(a), between each pair of CZs or dual CZs, both qubits are moved back to the storage zone and Pauli $X$ or $Z$ gates, respectively, are applied to each of them to cancel coherent dephasing from the movement and two-qubit gates. These baselines performed the best of those that we tried, disadvantaging any comparison between encoded and unencoded.
Results for both repeated logical CZ and repeated logical dual CZ, for up to 10 logical operations and 9 rounds of loss correction, are shown in Fig. 14, with detailed data in Table 6. Since the two physical baselines compile to the same circuit, we have combined them into a single baseline for comparison. The encoded repeated logical CZ experiments largely beat the corresponding baselines, including, for example, an encoded circuit consisting of 10 logical CZs with 9 loss-correction rounds outperforming an unencoded sequence of 10 CZs. Surprisingly, the repeated dual CZ experiment performs much worse for depths $8$ and $10$. Since it should be similar to repeated CZs, this could be from a mistake in scheduling the circuits, or possibly from machine parameter drift.
Figure 14: Error rates for $\llbracket 4,2,2\rrbracket$ code repeated logical CZ and repeated logical dual CZ experiments, and the corresponding physical baseline, conditioned on no initial loss. Error detection between logical gates (Fig. 13) permits mid-circuit correction of atom loss.
Table 6: Data for our $\llbracket 4,2,2\rrbracket$ repeated controlled-$Z$ and repeated dual controlled-$Z$ experiments. The encoded experiment with CZ or dual CZ depth $k+1$ has $k$ loss-correction rounds. It uses $5+2k$ physical qubits, with $5+8k$ physical CZ gates. A trial is prerejected if the cat state verification fails, and it is postrejected if an uncorrectable error is detected. Results for all available data, and for data in which there is no initial loss are reported separately. The unencoded baseline experiments, the same for CZ as for dual CZ, are on two qubits and include echos to cancel coherent errors. Trials with initial loss are prerejected, and those with circuit loss are postrejected. Restricting to trials with no initial loss has no effect on the baseline.
      All data
CZ or dual Loss-correction
CZ depth rounds Trials Prerejected Postrejected $\Pr[\text{accept}]$ Errors Error rate
Encoded CZs 2 1 3400 189 732 $73_{-2}^{+1}\%$ 34 $1.4^{+0.5}_{-0.4}\%$
4 3 4600 587 2030 $43(1)\%$ 43 $2.2^{+0.7}_{-0.6}\%$
6 5 3700 290 2408 $27(1)\%$ 41 $4(1)\%$
8 7 3700 302 2674 $20(1)\%$ 37 $5^{+2}_{-1}\%$
10 9 3700 318 2979 $10.9(1)\%$ 21 $5(2)\%$
Encoded dual CZs 2 1 5120 633 1051 $67(1)\%$ 53 $1.5(0.4)\%$
4 3 8640 1388 3340 $45(1)\%$ 104 $2.7(0.5)\%$
6 5 5120 1156 2899 $21(1)\%$ 69 $6^{+2}_{-1}\%$
8 7 5120 1270 3544 $7.7^{+0.8}_{-0.7}\%$ 63 $16^{+4}_{-3}\%$
10 9 5120 1640 3242 $4.7(6)\%$ 38 $16^{+5}_{-4}\%$
Baseline 2 4960 293 302 $88.0(9)\%$ 77 $1.8(4)\%$
4 4960 200 431 $87.3(9)\%$ 113 $2.6_{-0.4}^{+0.5}\%$
6 4960 202 566 $84.5(10)\%$ 186 $4.4_{-0.6}^{+0.7}\%$
8 4660 202 654 $82(1)\%$ 210 $5.5_{-0.7}^{+0.8}\%$
10 4960 284 860 $77(1)\%$ 374 $9.8_{-0.9}^{+1.0}\%$
      Trials with no initial loss
CZ or dual Loss-correction
CZ depth rounds Trials Prerejected Postrejected $\Pr[\text{accept}]$ Errors Error rate
Encoded CZs 2 1 3119 127 603 $77_{-2}^{+1}\%$ 21 $0.9^{+0.4}_{-0.3}\%$
4 3 3787 197 1692 $50(2)\%$ 28 $1.5^{+0.6}_{-0.5}\%$
6 5 3209 205 2048 $30(2)\%$ 32 $3(1)\%$
8 7 2890 188 2026 $23(2)\%$ 29 $4^{+2}_{-1}\%$
10 9 3004 210 2417 $13(1)\%$ 15 $4(2)\%$
Encoded dual CZs 2 1 4282 333 746 $75(1)\%$ 31 $1.0^{+0.4}_{-0.3}\%$
4 3 6732 854 2324 $53(1)\%$ 58 $1.6^{+0.5}_{-0.4}\%$
6 5 4064 793 2302 $24(1)\%$ 48 $5(1)\%$
8 7 3626 775 2500 $9.7^{+1.0}_{-0.9}\%$ 50 $14^{+4}_{-3}\%$
10 9 3312 940 2171 $6.1(8)\%$ 26 $13^{+5}_{-4}\%$
VI.2 Random circuits
As explained in Sec. III, the $\llbracket 4,2,2\rrbracket$ code allows a variety of other fault-tolerant logical gates. Each of $\{\mathrm{CNOT},\mathrm{CZ},\mathrm{SWAP},H^{\otimes 2}\}$ can be implemented by permutation or transversal gates. This set generates a group of size $36$.
To test the performance of these gates, and in the spirit of Ref. Got (16), we design experiments by preparing encoded ${|00\rangle}$ and then applying up to four elements of the group at random, with error detection between each pair of elements. The sequences chosen are, using CNOT = CX for brevity:
$0:$ $H^{\otimes 2}\,\mathrm{CX}$ (1)
$1:$ $I,\;\mathrm{CX}\,\mathrm{CZ}\,H^{\otimes 2}\,\mathrm{CX}$
$2:$ $H^{\otimes 2}\,\mathrm{SWAP}\,\mathrm{CX},\;\mathrm{CZ}\,H^{\otimes 2}\,\mathrm{SWAP},\;H^{\otimes 2}\,\mathrm{SWAP}\,\mathrm{CZ}$
$3:$ $\mathrm{CX}\,H^{\otimes 2}\,\mathrm{CX},\;H^{\otimes 2},\;\mathrm{SWAP}\,\mathrm{CX},\;\mathrm{CZ}\,\mathrm{CX}\,H^{\otimes 2}$
The encoding of sequence $j$ involves $j$ rounds of error detection. Note that the error detection circuits used here are those of Fig. 12(a), instead of those of Fig. 12(b); therefore they correct for lost qubits only at the end of the circuit. Future work may include executing these logical operation sequences with the circuits of Fig. 12(b).
The choice to error-detect between group elements is justified by imagining a computation with many [[4,2,2]] blocks. In that scenario, it is natural to write a circuit as alternating layers of CNOTs between blocks, and group elements within each block. It is then also natural to schedule error detection following each pair of layers in order to control the spread of errors among blocks.
The output of each random sequence is a stabilizer state with two generators. Each generator is measured to determine success of failure of the experiment. The encoded circuits can be directly compared to corresponding physical baselines that execute the same gate sequence. Results are shown in Fig. 15, with more details in Table 7. We find the encoded circuits consistently outperform the corresponding physical unencoded baselines.
Figure 15: Error rates for $\llbracket 4,2,2\rrbracket$ code random logical circuits, conditioned on no initial loss. Sequence labels are from the sequences given in Eq (1).
Table 7: Data for our $\llbracket 4,2,2\rrbracket$ random circuit experiments. The sequences of operations are specified in Eq. (1). The encoding of sequence $k$ includes $k$ rounds of error detection, $5+2k$ physical qubits and $5+8k$ physical CZ gates. For sequences 0, 1, and 2, the two stabilizer generators of the output each have weight one and are measured simultaneously in each experiment. An error is reported if either of the two generators have incorrect eigenvalue. The stabilizer generators of sequence 3 are $\{XZ,ZX\}$. The encoded circuit measures both of these generators in the same experiment via transversal $Y$-basis measurements. The physical baselines require separate experiments for each generator. Results for the two generators are therefore reported separately. All data is reported. Restricting to trials with no initial loss has no effect on the baseline.
    All data
Sequence Trials Pre. Post. $\Pr[\text{accept}]$ Errors Error rate
Encoded 0 1200 66 $42$ $91(2)\%$ 6 $0.6^{+0.6}_{-0.3}\%$
1 2800 119 $467$ $79_{-2}^{+1}\%$ 5 $0.2^{+0.4}_{-0.1}\%$
2 2500 132 $737$ $65(2)\%$ 9 $0.6^{+0.4}_{-0.3}\%$
$XZ$ 3 7200 595 $3292$ $46(1)\%$ 15 $0.5^{+0.3}_{-0.2}\%$
$ZX$ 3 7200 595 $3292$ $46(1)\%$ 24 $0.7(3)\%$
Baseline 0 1600 52 $51$ $94(1)\%$ 20 $1.3^{+0.7}_{-0.5}\%$
1 1600 40 $55$ $94(1)\%$ 13 $0.9^{+0.6}_{-0.4}\%$
2 1600 66 $47$ $93(1)\%$ 23 $1.6^{+0.7}_{-0.5}\%$
$XZ$ 3 1800 87 $108$ $89(1)\%$ 50 $3.1^{+0.9}_{-0.8}\%$
$ZX$ 3 1800 108 $134$ $87(2)\%$ 40 $2.6^{+0.9}_{-0.7}\%$
Trials with no initial loss
Sequence Trials Pre. Post. $\Pr[\text{accept}]$ Errors Error rate
Encoded 0 1104 45 $35$ $93_{-2}^{+1}\%$ 3 $0.3^{+0.5}_{-0.2}\%$
1 2645 95 $398$ $81_{-2}^{+1}\%$ 3 $0.1^{+0.2}_{-0.1}\%$
2 2289 105 $605$ $69(2)\%$ 9 $0.6^{+0.5}_{-0.3}\%$
$XZ$ 3 6315 425 $2702$ $50(1)\%$ 15 $0.5^{+0.3}_{-0.2}\%$
$ZX$ 3 6315 425 $2702$ $50(1)\%$ 24 $0.8(3)\%$
VII Repeated error correction with the distance-three Bacon-Shor code
Running deep logical circuits in a neutral atom quantum computer will ultimately require repeated correction of both errors within the qubit subspace and atom loss from erasure conversion. In order to demonstrate this, we choose the distance-three $\llbracket 9,1,3\rrbracket$ Bacon-Shor code. This code has previously been implemented on trapped-ion systems EDN+ (21); HBC (24). A code distance greater than two is needed to correct errors other than loss.
To our knowledge, the present work is the first demonstration of repeated error correction in neutral atoms. We employ a scheme similar to that of Steane and Knill Ste (97); Kni (05); PSR+ (24), leveraging its ability to correct atom loss; see Fig. 16. In this work, the ancilla qubits are held until readout at the end of the circuit, so each round of error correction requires the use of additional qubits. However, in principle, this scheme can be combined with mid-circuit measurement and reset where readout and loss detection, combined with mid-circuit reloading, will enable arbitrary-depth circuits.
[…]
After one round of error correction, we find a logical error rate of $4.9^{+1.0}_{-0.9}\%$ and after two rounds the error rate is $8^{+2}_{-1}\%$. Table 8 gives more details. While these results suggest the opportunity to further improve physical error rates, they also serve as a demonstration of the mechanism to correct for both errors and loss. This can be seen in the analysis of the successful trials with one round of error correction, where on average $0.20$ erasures and at least $0.14$ errors were corrected in the readout of the final data block, whereas $0.40$ erasures and $0.50$ errors were corrected in the previous round of $X$ and $Z$ error correction, a higher average number of corrections. In the successful trials with two rounds of error correction, on average $0.22$ erasures and at least $0.13$ errors were corrected in the readout of the final data block (consistent with the one-round data), while $0.95$ erasures and $0.88$ errors were corrected in the previous error-correction rounds.
In order to compare encoded and physical results, we must choose a meaningful physical baseline. From the standpoint of the encoded data, the logical circuit is a quantum memory, applying the logical identity operation. For a meaningful comparison, we show results for repeated teleportation of physical qubit states, following the pattern of repeated teleportation of the logical data in Fig. 16(a).
Table 8: Data for $\llbracket 9,1,3\rrbracket$ Bacon-Shor code repeated error correction (EC) experiments. The data for the encoded circuits includes all trials, including those with initial loss, and there is no preselection. When decoding logical $X$ ($Z$), a trial is postrejected if there are losses in two or more rows (columns) of the $3\times 3$ code, or if one row is lost and the other two disagree. For the unencoded baseline, a trial is prerejected if there are any initial losses and a surviving trial is postrejected if there are any circuit losses.
  Physical Physical
Experiment qubits CZs Trials Prerejected Postrejected $\Pr[\text{accept}]$ Errors Error rate
Encoded 1 EC round 27 36 2400 541 $77(2)\%$ 91 $4.9^{+1.0}_{-0.9}\%$
2 EC rounds 45 66 2400 1054 $56(2)\%$ 114 $8^{+2}_{-1}\%$
Baseline 2 1Q teleportations 3 2 $640$ 31 47 $88^{+2}_{-3}\%$ 7 $1.3^{+1.2}_{-0.7}\%$
4 1Q teleportations 5 4 $640$ 50 82 $79(3)\%$ 7 $1.4^{+1.3}_{-0.8}\%$
VIII Outlook
In this work, we have created large entangled logical states, implemented computational algorithms on encoded qubits, demonstrated qubit loss correction, and implemented a code with the capacity for repeated loss and error correction. Altogether, these results highlight the unique features and promise of neutral atoms as a platform for reliable fault-tolerant quantum computing.
The large encoded cat states and the Bernstein-Vazirani algorithms are enabled by the combination of scalability, high-fidelity operations inherent to neutral atoms, and all-to-all connectivity among the qubits. For larger-scale logical computation, this connectivity paired with high qubit count will enable the use of nonlocal error-correcting codes with high encoding rates and low physical qubit overhead XZZ+ (24).
All of the encoded experiments, particularly those with repeated error detection, highlight the usefulness of loss-to-erasure conversion in the context of encoded qubits. They demonstrate that logical circuits can operate effectively with atom loss, which will be a key feature of large-scale neutral atom quantum computers.
This work and similar work with other neutral atom platforms BEG+ (24), as well as recent work with superconducting qubits GC (24) and trapped ion qubits PSR+ (24); RAC+ (24), highlight a critical advance in the field of quantum computation: the transition from scientific results computed with physical qubits to computational advances achieved with logical qubits. After a relatively short time the neutral atom platform has shown significant advancement from its inception as an experimental quantum computing platform. In the near future we expect advances in both 2Q gate fidelities as more iterations are performed to reduce gate laser noise as well as significant scale up towards 10 0001000010\,00010 000 qubits allowing for larger distance codes, more performant logical qubits, and more impressive demonstrations of fault-tolerant, universal quantum computation.
The combination of a hardware-optimized qubit virtualization system, complete with high-rate and high-efficiency quantum error correction, with a programmable neutral atom quantum processor, inclusive of mid-circuit measurement NC+ (23), continuous atom reloading NKC+ (24), and increased two-qubit gate fidelity, is a promising path to achieving large-depth, logical computation for scientific quantum advantage.
Acknowledgments
We thank Zulfi Alam, David Bohn, Jeongwan Haah, Chetan Nayak, Pasi Kostamo, Vadym Kliuchnikov, Dennis Tom, Matthias Troyer, and Jason Zander for their support and feedback during this project.

### v3

@@
-Logical computation demonstrated with a neutral atom quantum processor
+Fault-tolerant quantum computation with a neutral atom processor
 Abstract
-Transitioning from quantum computation on physical qubits to quantum computation on encoded, logical qubits can improve the error rate of operations, and will be essential for realizing valuable quantum computational advantages.
-Using a neutral atom quantum processor with $256$ qubits, each an individual Ytterbium atom, we demonstrate the entanglement of $24$ logical qubits using the distance-two $\llbracket 4,2,2\rrbracket$ code, simultaneously detecting errors and correcting for lost qubits.
-We also implement the Bernstein-Vazirani algorithm with up to $28$ logical qubits encoded in the $\llbracket 4,1,2\rrbracket$ code, showing better-than-physical error rates.
-We demonstrate fault-tolerant quantum computation in our approach, guided by the proposal of Gottesman (2016), by performing repeated loss correction for both structured and random circuits encoded in the $\llbracket 4,2,2\rrbracket$ code.
-Finally, since distance-two codes can correct qubit loss, but not other errors, we show repeated loss and error correction using the distance-three $\llbracket 9,1,3\rrbracket$ Bacon-Shor code.
+Quantum computing experiments are transitioning from running on physical qubits to using encoded, logical qubits.
+Fault-tolerant computation can identify and correct errors, and has the potential to enable the dramatically reduced logical error rates required for valuable algorithms.
+However, it requires flexible control of high-fidelity operations performed on large numbers of qubits.
+We demonstrate fault-tolerant quantum computation on a quantum processor with $256$ qubits, each an individual neutral Ytterbium atom.
+The operations are designed so that key error sources convert to atom loss, which can be detected by imaging.
+Full connectivity is enabled by atom movement.
+We demonstrate the entanglement of $24$ logical qubits encoded into $48$ atoms, at once catching errors and correcting for, on average 1.8, lost atoms.
+We also implement the Bernstein-Vazirani algorithm with up to $28$ logical qubits encoded into $112$ atoms, showing better-than-physical error rates.
+In both cases, “erasure conversion,” changing errors into a form that can be detected independently from qubit state, improves circuit performance.
 These results begin to clear a path for achieving scientific quantum advantage with a programmable neutral atom quantum processor.
-I Introduction
-Quantum computers can enable faster, more accurate solutions than classical computers.
-However, the biggest challenge to scaling up quantum computers is errors, or “noise.” Physical qubits and their operation are inherently noisy.
-State-of-the-art two-qubit gate error rates depend on the qubit implementation, and in some flagship quantum processors range, e.g., from $p=0.1\%$ for an ion trap processor MBA+ (23) to an average of $p=0.4\%$ for a superconducting qubit processor GC (24).
-If a circuit cannot tolerate errors, it can only run for roughly $1/p$ steps, or fewer, which is not nearly enough to achieve large-scale, advantageous computation.
-Even the best physical qubits are insufficiently reliable, without error correction, to run even moderate-size quantum algorithms.
-Indeed, obtaining a quantum advantage in scientific applications will likely require at least $100$ qubits with error rate $p$ at most $10^{-6}$, while a quantum advantage for industrial-scale problems may require upwards of $1000$ qubits with $p\sim 10^{-12}$ BMT+ (22).
-Such error rates will not be achievable at the physical level.
-Fortunately, encoding qubits in quantum error-correcting codes, thereby creating “logical” qubits, should allow dramatic reductions in effective error rates.
-The operations must also be encoded, according to fault-tolerant constructions AGP (06).
-Thus, by using quantum error correction and fault tolerance, noise in quantum hardware can be addressed through software, albeit with a cost in physical qubit overhead.
-In this work, we present a step toward achieving scientific quantum advantage by demonstrating reliable quantum computation using logical qubits that outperforms corresponding physical-qubit based computation.
-We present our results on a neutral atom quantum processor.
-Neutral atom quantum architectures recently have been in rapid development.
-Previously, neutral atom architectures were predominantly analog in nature.
-Recent progress has enabled neutral atom approaches to transition to digital, gate-based computation EBK+ (23); SST+ (23); MLP+ (23); RCC+ (24); FTS+ (24); MSS+ (24).
-The promise of neutral atoms as qubits is their configurability, both in count and connectivity, while also maintaining low error rates.
-The neutral atom quantum processor we use in this work has up to 256 qubits (Ytterbium atoms), with all-to-all connectivity enabled by atom movement, along with high two-qubit physical gate fidelity and atom loss detection.
-This allows for flexible computation, as well as the ability to both use and optimize over a wide range of quantum error-correcting codes, including those requiring non-local connectivity, for the underlying neutral atom architecture.
-We refer to the system for detecting and correcting errors, also during computation, as the qubit virtualization system, as encoded, logical qubits are virtual qubits made up of some number of physical qubits.
-Through co-optimization of fault-tolerant quantum error correction protocols within the qubit virtualization system and the neutral atom quantum processor design, we demonstrate the creation of logical qubits, their logical entanglement, and logical computation, while correcting for lost atoms.
-Figure 1: Summary of results for the $24$-qubit cat state.
+Quantum computers are potentially powerful, but are challenging to control precisely.
+State-of-the-art two-qubit gate error rates in some flagship quantum processors range from $0.1\%$ for a trapped-ion device MBA+ (23) to an average of $0.4\%$ for a superconducting qubit device GC (24).
+Moderate-size quantum algorithms require much lower error rates, below $10^{-6}$ BMT+ (22).
+This precision can likely only be achieved using quantum error-correcting codes to store and fault-tolerantly operate on “logical” qubits AGP (06).
+Here, we advance the shift from physical to logical qubits, by demonstrating fault-tolerant, encoded computations on logical qubits.
+The encoded circuits outperform unencoded circuits on physical qubits.
+Our results are achieved using a neutral atom quantum processor with $256$ qubits, Ytterbium atoms (Fig.
+1), designed to enable loss conversion and detection for improved logical performance.
+Neutral atom devices have been mainly analog, but recent progress has enabled digital, gate-based computation EBK+ (23); SST+ (23); MLP+ (23); RCC+ (24); FTS+ (24); MSS+ (24).
+The processor here has low gate error, and atom movement enables all-to-all connectivity for flexible computation and logical operation design.
+The physical two-qubit gates are designed so that key error sources convert to atom loss, which imaging has been designed to detect.
+With unencoded circuits, data with lost atoms can be discarded, and using this selection we prepare an entangled $40$-qubit “cat” state (Fig.
+2).
+Loss conversion is even more advantageous operating on encoded data, though; a distance-two code can detect errors and correct for lost atoms.
+We demonstrate the creation and entanglement of logical qubits, and logical computation.
+We prepare a cat state on $24$ logical qubits, encoded in $48$ atoms (Fig.
+3).
+We then demonstrate a quantum algorithm on logical qubits.
+We fault-tolerantly run the Bernstein-Vazirani algorithm BV (97); CEMM (98), well known for showing the higher efficiency of quantum computers versus classical, on as many as $28$ logical qubits, encoded in $112$ atoms (Fig.
+4), with better-than-physical results.
+This illustrates the scale available in neutral atom processors, and their promise as reliable, fault-tolerant quantum processors.
+Figure 1: Experimental architecture.
+a, b, 171Yb atoms, each giving a physical qubit in the ground-state nuclear spin, can occupy tweezer traps within either a register array or an interaction zone (IZ).
+Up to 8 pairs of atoms interact at a time in the IZ, driven by two-qubit (2Q) gate lasers to implement controlled-phase (CZ) gates.
+These lasers perform coherent sequential excitation to a metastable clock state (3P0) and the Rydberg state (653S1).
+Arbitrary single-qubit (1Q) operations can be applied in parallel to the atoms within a register row using Raman transitions detuned from the 3P1 manifold.
+Moving atoms between the zones enables arbitrary 2Q gate connectivity.
+c, Circuit-running sequence.
+Dissipative operations (left column) are performed with atoms held in a cavity-enhanced optical lattice.
+Three images (orange boxes) allow us to determine initial occupancy of register sites, as well as the final state and presence of atoms.
+Coherent operations (right column) are performed with atoms held in tweezers.
+Figure 2: Fidelity of prepared cat states versus qubit number $n$.
+Red points give fidelities for trials with no lost atoms.
+Cat states with up to $40$ qubits exceed the entanglement threshold.
+Gray points include trials in which all atoms are present initially but atoms lost during the circuit are replaced by the maximally mixed state, and black points include all trials.
+Figure 3: Encoded $24$-qubit cat state results.
 Shown is the error for measurements in the $X$ basis, the error for measurements in the $Z$ basis, and their sum, for three cases.
-A total less than $50\%$ implies entanglement.
-The baseline (left) is an unencoded $24$-qubit cat state; the error bars are large because we have conservatively chosen only the best run instead of averaging all baseline runs, which would further increase the separation between logical and physical performance.
-See Fig.
-7 for more details.
-The encoded bars are based on different ways of interpreting the same data for $\llbracket 4,2,2\rrbracket$-encoded cat states.
-In one case (middle), we reject trials with a detected error, but attempt to correct for lost qubits.
-In the other case (right), we reject trials with any lost qubit or detected error.
-Figure 2: Summary of results for the Bernstein-Vazirani algorithm with an $n$-bit secret string.
+The baseline is an unencoded $24$-qubit cat state, based on trials with no lost atoms.
+For the encoded cat state data, in one case we reject trials with a detected error, but try to correct for qubits lost during the circuit.
+In the other case, we reject trials with any lost qubit or detected error.
+Figure 4: Results for the Bernstein-Vazirani algorithm with an $n$-bit secret string.
 The string is always $1^{n}$, as the corresponding oracle is the most difficult to implement.
@@
 The lighter points are based only on those trials with no initial atom loss.
-Key results.
-We demonstrate a variety of reliable quantum computations that outperform physical qubit computation using logical qubits co-designed with a programmable neutral atom quantum processor.
-First, we show entanglement of $24$ logical qubits in a “cat” state (Fig.
-1).
-This is the largest encoded cat state on record to date, illustrating the scale available in neutral atom processors.
-We then demonstrate quantum algorithms encoded into logical qubit computations.
-We implement the Bernstein-Vazirani algorithm BV (97); CEMM (98), a quantum algorithm well known for showing the higher efficiency of quantum computers versus classical.
-Our implementations involve as many as $28$ logical qubits, encoded in $112$ physical qubits (Fig.
-2).
-Next, we demonstrate fault tolerance guided by a proposal of Gottesman Got (16).
-These experiments incorporate repeated error detection and, notably, repeated loss correction, a critical feature for scaling neutral atom processors.
-Each of the above experiments are encoded to distance two; they are capable of detecting phase and bit-flip errors and correcting atom loss.
-Finally, we demonstrate correcting both loss and errors, with repeated error correction in the distance-three Bacon-Shor code Bac (06).
-Related work.
-Creating and verifying large cat states, also known as GHZ states GHZ (89), is a common benchmark for measuring hardware progress Kre (24).
-The quality of an experiment is typically expressed by the fidelity with the ideal cat state, and where a fidelity greater than $1/2$ is the success criterion.
-For example, Moses et al.
-MBA+ (23) prepared the $20$-qubit cat state with an $86\%$ fidelity, and the $32$-qubit cat state with an $82\%$ fidelity, on the Quantinuum H2 trapped-ion processor.
-Using superconducting qubits, Bao et al.
-BXS+ (24) prepared a $60$-qubit cat state with a $59\%$ fidelity.
-For encoded qubits, the published record to date is an encoded $12$-qubit cat state with at least a $99.8\%$ fidelity, also on H2 RAC+ (24).
+Creating and verifying large cat or GHZ states GHZ (89) is a common benchmark for hardware progress Kre (24).
+The current record is a $60$-qubit cat state, prepared with $59\%$ fidelity using superconducting qubits BXS+ (24).
+A $20$-qubit cat state with $54\%$ fidelity, after adjusting for measurement error, has been prepared on a neutral atom system OLK+ (19).
+The $62^{+5}_{-6}\%$ fidelity of our $40$-qubit cat state is not adjusted for measurement errors.
+Table 1 summarizes results with encoded cat states.
+A cat state on $12$ logical qubits has been prepared with $99.8\%$ fidelity RAC+ (24), on a trapped-ion processor MBA+ (23).
 On a neutral atom system, Bluvstein et al.
-BEG+ (24) have prepared a four-qubit encoded cat state with a fidelity of $72(2)\%$ using error correction, ranging to $99.85^{+0.1}_{-1.0}\%$ using full error detection and therefore with much lower acceptance probability.
-Table 1 summarizes results with encoded cat states.
+BEG+ (24) have prepared a four-qubit encoded cat state with fidelity $72(2)\%$ using error correction, ranging to $99.85^{+0.1}_{-1.0}\%$ using full error detection and with a lower acceptance rate.
+Our $24$-qubit encoded cat state is the largest published to date.
 Table 1: Experiments preparing encoded cat states.
   Logical
-Reference qubits Code Fidelity
-HDHL (24) 4 $\llbracket 25,4,3\rrbracket$ $99.5^{+0.2}_{-0.4}\%$ to $99.7^{+0.2}_{-0.3}\%$
+Reference qubits Code Total error $p_{X}+p_{Z}$
+HDHL (24) 4 $\llbracket 25,4,3\rrbracket$ $0.5^{+0.4}_{-0.2}\%$
 BEG+ (24) 4 $\llbracket 7,1,3\rrbracket$
 […]
-72(2)error correction
-$99.85^{+0.1}_{-1.0}\%$ error detection
-RAC+ (24) 12 $\llbracket 16,4,4\rrbracket$ $99.82^{+0.12}_{-0.4}\%$ to $99.90^{+0.1}_{-0.3}\%$
+28(2)error correction
+$0.15^{+1.0}_{-0.1}\%$ error detection
+RAC+ (24) 12 $\llbracket 16,4,4\rrbracket$ $0.18^{+0.4}_{-0.12}\%$
 This work 24 $\llbracket 4,2,2\rrbracket$
 […]
-$73.4^{+3.1}_{-3.3}\%$ to $85.6^{+2.3}_{-2.5}\%$
-error detection
+$26.6^{+3.3}_{-3.1}\%$ error detection
 and loss correction
-$89.8^{+4.7}_{-6.6}\%$ to $91.5^{+4.3}_{-6.2}\%$
-error and loss detection
+$10.2^{+6.6}_{-4.7}\%$ error and
+loss detection
 The Bernstein-Vazirani algorithm for finding a secret string is another benchmark for quantum processors LGM+ (24).
@@
 Some test identification of the secret $s=1^{n}$, since it is hardest to implement; the corresponding oracle requires the most two-qubit gates.
-Others test against random strings $s$ or average over different Hamming weights, and this is much easier, particularly on ion traps, comparable to finding $s=1^{n/2}$.
-The success probabilities also vary considerably, e.g., from an exponentially small value for random $s$ with length $26$ PL (23), to $70\%$ for random $s$ with length $20$ LJV+ (23).
-A critical capability for enabling more complex, reliable computation in a quantum processor is repeated detection and correction of errors.
-The latest generations of quantum hardware have enabled a variety of such demonstrations within the past few years.
-Repeated error detection with distance-two codes has been demonstrated on ion traps, including three rounds on eight logical qubits SBA (24), and 15 rounds on four logical qubits YDKR (24).
-Demonstrations based on Gottesman’s protocol Got (16) for fault tolerance with the $\llbracket 4,2,2\rrbracket$ code, but without including repeated error detection, include WWJ+ (18); HF (19); Vui (18); LGL+ (17); CCNH (21).
-Repeated error correction, with higher-distance codes, has also been demonstrated.
-For single logical qubits, demonstrations include those on trapped-ion processors RBL+ (21); PBP+ (24) and superconducting processors KLR+ (22); Goo (23); GC (24); PNH+ (24) each with varying numbers of rounds, e.g., up to nearly $36$ rounds of error correction ($250$ syndrome extraction cycles) on the $\llbracket 49,1,7\rrbracket$ surface code, with a $0.143(3)\%$ logical error rate per cycle GC (24).
-Repeated error correction on multiple qubits has been implemented, including three rounds on two logical qubits PSR+ (24), and five rounds on eight logical qubits RAC+ (24), each while beating error rates of physical baselines.
-II Physical machine description
-Figure 3: Diagram of experimental architecture.
-171Yb atoms, which each encode a physical qubit in their ground-state nuclear spin, can occupy tweezer traps within either a register array or an interaction zone (IZ).
-Up to 8 pairs of atoms interact at a time in the IZ, driven by two-qubit (2Q) gate lasers to implement controlled phase gates on the atoms in each pair.
-Arbitrary single-qubit (1Q) operations can be applied in parallel to the atoms within a row of the register; here one atom is missing.
-Each atom has a home register location, indicated by dashed lines, where it is originally loaded and where 1Q gates can be applied.
-Movement of atoms between the two zones enables arbitrary connectivity for 2Q gates.
-Up to eight atoms from the same row can move in parallel, but their paths cannot cross.
-All results presented in this work were obtained on a quantum processor based on reconfigurable arrays of neutral 171Yb atoms, depicted in Fig.
-3, with the qubits encoded in the ground-nuclear spin states (1S0, $m_{F}=\pm 1/2$).
-Aspects of the system have previously been described in BBB+ (22); NC+ (23); NKC+ (24), as well as an accompanying publication MSS+ (24).
-Further system and benchmarking details will follow in a separate manuscript, in preparation.
-By moving atoms between arrays of static optical tweezer traps using mobile tweezers, this platform enables arbitrary connectivity between qubits, for a large numbers of qubits.
-Single-qubit operations and qubit storage are performed in a register array, which is loaded in an iterative manner from a repeatedly filled reservoir NKC+ (24).
-Two-qubit gates are performed in a separate interaction zone (IZ).
-For this work, up to 256 atoms are available for computation, and up to eight atoms within the same row may be moved between the register and IZ at a time, subject to constraints associated with intermodulation in the acousto-optic deflectors used to generate the moving tweezers.
-Typical atom loss probabilities per move are below 0.1%.
-State-resolved, low-loss readout is performed by transferring atoms from the tweezer arrays into traps formed by cavity-enhanced optical lattices, which allows for state and atom detection with error and loss rates at the few parts-per-thousand level (see NKC+ (24) for a detailed characterization of state preparation and imaging).
-Importantly, the ability to perform state-selective, low-loss imaging allows us to distinguish between different types of error.
-An image performed prior to circuit execution allows the identification of sites successfully occupied by atoms, while a pair of images directly after circuit execution and after optical pumping within the qubit subspace allow determination of both the presence and state of the atoms.
-In the encoded Bernstein-Vazirani experiment with $n=7$, for example, on average $0.52(7)\%$ of the $32$ atoms were not loaded, and $1054$ of $1200$ shots, or $85(2)\%$, had no initial loss.
-The fraction of qubits lost during the circuit averaged $2.3^{+0.2}_{-0.1}\%$.
-Initial loss varies, and more complicated circuits would have a higher circuit loss rate.
-Single-qubit operations are performed in parallel on rows of atoms within the register, with the pulse area, phase, and presence of the gates controlled individually on each site using an architecture similar to that described in Ref.
-BBB+ (22).
+Others test against random strings $s$, and this is much easier, comparable to finding $s=1^{n/2}$.
+The success probabilities also vary considerably, e.g., from $0.2\%$ for $s=1^{26}$ PL (23), to $70\%$ for random $s$ with $n=20$ LJV+ (23).
+Our fault-tolerant, encoded implementation of the Bernstein-Vazirani algorithm is the first to date, and has the largest length, $n=27$.
+On $s=1^{27}$, the $91.4(3)\%$ success probability for the encoded circuit beats the unencoded algorithm.
+Neutral atom quantum processor
+The quantum processor used in this work is based on reconfigurable arrays of neutral 171Yb atoms, depicted in Fig.
+1a,b, with the qubits encoded in the ground-nuclear spin states (1S0, $m_{F}=\pm 1/2$).
+Aspects of the system have been described in previous publications BBB+ (22); NC+ (23); NKC+ (24); MSS+ (24), and further details can be found in the methods section.
+By moving atoms between arrays of static optical tweezer traps using mobile tweezers, this platform enables arbitrary connectivity between large numbers of qubits.
+Figure 1c shows a typical sequence for circuit execution.
+State preparation consists of iterative loading of the relevant sites of the register array, as described in NKC+ (24), followed by an image to determine successful filling of sites, cooling, and initialization of the ${|0\rangle}$ state through optical pumping.
+State-resolved, low-loss readout as well as cooling and state initialization are performed by transferring atoms from the tweezer arrays into traps formed by cavity-enhanced optical lattices, which allows for state and atom detection with error and loss rates at the part-per-thousand level (see methods).
+All other operations are peformed with the atoms held in the tweezer arrays.
+For this work, up to 256 atoms are available for computation.
+During circuit execution, single-qubit operations and qubit storage are performed in a register array.
+Two-qubit (2Q) gates are performed on up to eight qubit pairs in parallel within a separate interaction zone (IZ), with connectivity determined by the placement of atoms within the IZ.
+2Q gates are performed using state-selective sequential two-photon excitation to a high-lying Rydberg state in a blockaded regime, resulting in controlled-phase (CZ) gates between pairs of qubits.
+Importantly, key sources of 2Q gate errors are converted to loss, which can be detected independently from the qubit state.
+Population errors in the sequential excitation used in the gates lead either directly to loss, or to population of metastable states that are not trapped in register traps.
+Unlike most systems based on individually controlled neutral atoms, our ability to perform nondestructive state-selective imaging NC+ (23) allows us to perform a pair of images that unambiguously distinguish between three possibilities for the atomic state: ${|0\rangle}$, ${|1\rangle}$, and lost.
+Table 2 presents a summary of important error rates, with benchmark protocols described in the methods.
+Table 2: Randomized benchmark (RB) infidelities, conditioned on an atom (1Q) or pair of atoms (2Q) staying in the qubit subspace, and probability of losing one or both atoms.
+  Benchmark protocol Infidelity Loss + Leakage
+2Q Static IRB 0.56(15)0.27(6)%
+2Q Echoed RB with moves 0.39(10)0.75(10)%
+2Q Echoed RB without moves 0.35(5)0.24(6)%
+Cat state generation
+We prepare non-encoded cat states using modified versions of the circuits from CFG+ (19).
+We generate the “antiferromagnetic” cat state $\frac{1}{\sqrt{2}}({|0101\ldots\rangle}+{|1010\ldots\rangle})$, instead of the “ferromagnetic” state $\frac{1}{\sqrt{2}}({|0^{n}\rangle}+{|1^{n}\rangle})$, since it is less sensitive to correlated phase errors.
+Figure 2 plots the fidelity to the ideal cat state, computed as $1-\tfrac{1}{2}(Z$ basis error rate)()-() - (average error rate over $n$ bases orthogonal to $z)$ GLGP (07).
+A fidelity above 50% certifies $n$-partite entanglement GS (10).
+Restricting the analysis to only those trials where all atoms survived, the cat states exceed the entanglement threshold up to $n=40$.
+To our knowledge, this is the largest cat state demonstrated with neutral atoms.
+We can also estimate the fidelity including all trials, or all trials with no initial loss, in the analysis generously treating lost qubits as replaced by the maximally mixed state.
+This gives a higher acceptance rate, but lower fidelity, showing the advantage of loss detection.
+Encoded $24$-qubit cat state
+We prepare a cat state on $24$ logical qubits, the largest published to date.
+The logical qubits are encoded with the $\llbracket 4,2,2\rrbracket$ code into $48$ physical qubits.
+Starting with a logical state-preparation circuit, we encode the circuit, then schedule the gates and movements (Fig.
+5).
+Figure 5: a, Logical circuit to prepare a $24$-qubit cat state.
+The qubits are arranged in $12$ pairs that will each be encoded into a $\llbracket 4,2,2\rrbracket$ code block.
+b, The encoded circuit involves $59$ atoms and $101$ physical CZ gates.
+All but the first code block use extra ancilla qubits, shown in gray, to prepare encoded ${|00\rangle}$ fault tolerantly; each ancilla should be measured to be 00.
+In the diagram, a blue wire is a qubit in the Hadamard basis.
+Not shown are the final transversal $X$ or $Z$ measurements.
+c, Schedule for the atom movements and gates.
+Our experiments, like HDHL (24); RAC+ (24), measure $X$ basis and $Z$ basis error rates, $p_{X}$ and $p_{Z}$.
+In the unencoded case, the fidelity to the ideal cat state is at least $1-p_{X}-p_{Z}$.
+We find that for the encoded cat state, using error detection and loss correction, $p_{X}+p_{Z}=26.6^{+3.3}_{-3.1}\%$, while for the unencoded baseline $p_{X}+p_{Z}=42.0^{+11.3}_{-10.1}\%$.
+The results are summarized in Fig.
+3.
+The baseline error bars are large because we report the best single run (different from in Fig.
+2), which disadvantages the encoded cases.
+The encoding is beneficial, but affects $p_{X}$ and $p_{Z}$ differently.
+How useful is loss conversion, loss measurement, and loss correction for this encoded state?
+The average loading loss rate was $1.04(2)\%$, and $57.2(6)\%$ of trials had no initially lost atoms.
+Among trials without initial losses, those that decoded correctly had an average of 2.0 and 1.6 circuit losses, for $X$ and $Z$ basis measurements, respectively.
+The code successfully corrects for these lost qubits.
+Still, circuit loss is a major cause of logical errors.
+Those shots that decoded to a logical error had respective averages of 2.4 and 2.7 losses.
+In the $Z$ basis, there are only two valid decodings, $0^{24}$ and $1^{24}$, and we can usually identify the most likely cause of a logical error.
+For example, a decoding of $1111(0111)^{5}$ indicates that most probably the third block had a logical $XI$ error that propagated to the subsequent code blocks.
+The frequencies of the most likely causes of the logical error are: $26\%$ two circuit losses, $54\%$ one loss and one $X$ error, $9\%$ two $X$ errors, $11\%$ three or more losses or errors.
+Figure 6 shows how conditioning on at most $k$ lost qubits, for $k=0,1,\ldots$, allows an interpolation between the different encoded results in Fig.
+3; in particular, conditioning on no loss reduces the total error from $29.5(3.0)\%$ to $10.2^{+6.6}_{-4.7}\%$.
+Finally, Fig.
+7 shows all the encoded cat state runs, plotted conditioned on no initial atom loss.
+The spread and correlations in the figure suggest some machine parameter drift.
+Figure 6: Encoded $24$-qubit cat state data, selecting only trials with specified loss criteria.
+Trying to correct more loss generally leads to a higher acceptance rate, but also a higher error rate.
+In each trial, an atom may be missing before the circuit begins or it may be lost during the circuit.
+The red points show the total $X$ basis plus $Z$ basis error rate, plotted against the average of the $X$ and $Z$ acceptance rates, based on either all 29 7602976029\,76029 760 trials, only those 17 0081700817\,00817 008 trials with no initial loss, or only those $858$ trials with no loss at all.
+The blue points select only trials in which the total number of lost qubits is at most a threshold.
+Figure 7: Encoded $24$-qubit cat state runs.
+Each point gives the acceptance probability and logical error rate for a run of, on average, $183$ trials with no initial loss, blue for the $X$ basis and red for $Z$.
+For clarity, only some error bars are shown.
+The large squares are the full averages, as reported in Fig.
+3.
+The negative correlation between acceptance probability and logical error rate evidences some parameter drift between runs, despite automatic recalibrations.
+Encoded Bernstein-Vazirani algorithm
+Computing with logical qubits is the next step after encoded state preparation.
+For a secret string $s\in\{0,1\}^{n}$, consider a black-box oracle that on query $x\in\{0,1\}^{n}$ returns $s\cdot x=s_{1}x_{1}\oplus\cdots\oplus s_{n}x_{n}$.
+Classically, $n$ oracle queries are required to learn $s$.
+The Bernstein-Vazirani (BV) algorithm BV (97); CEMM (98) makes a single call to the corresponding quantum oracle $U_{s}$, $U_{s}{|b,x\rangle}={|b\oplus(s\cdot x),x\rangle}$, to learn $s$ via $H^{n+1}U_{s}H^{n+1}{|1,0^{n}\rangle}={|1,s\rangle}$.
+This algorithm was designed to prove an oracle separation between the classical randomized and quantum polynomial-time complexity classes $\mathsf{BPP}$ and $\mathsf{BQP}$.
+[…]
+Figure 8: The Bernstein-Vazirani algorithm for a secret string $s\in\{0,1\}^{n}$ uses $n+1$ logical qubits.
+a, In the most challenging case, $s=1^{n}$ and there are $n$ logical CNOT gates.
+b, Encoded into the $\llbracket 4,1,2\rrbracket$ code, there are $4(n+1)$ atoms and $26n+$ 2 CZ gates.
+c, The gates are scheduled so that the first block, once initialized, stays in the interaction zone.
+As transversal CZ gates are applied into block $j$, block $j+1$ is initialized.
+The unencoded algorithm accepts if no measured atom is lost, and is successful if the measured string is $1^{n}$.
+The encoded algorithm tolerates losses, but rejects if any of the last $n$ code blocks is not decodable.
+Figure 4 summarizes the results. The encoded algorithms have at least a 1.4×1.4$\times$1.4 × lower error rate than the unencoded baseline, albeit with a lower acceptance rate.
+Computing the success probability conditioned on acceptance makes sense if we treat the BV algorithm as a generic quantum algorithm.
+However, it was originally proposed as a quantum query algorithm, in which minimizing the number of oracle queries was key.
+In this light, we can ask whether with only one query (trial), the encoded circuit beats the unencoded one.
+When the algorithms cannot condition on acceptance, it requires guessing $s$.
+With a lost unencoded qubit or an undecodable code block the probability of guessing the corresponding bit of $s$ is $1/2$, and the expected Hamming distance from $s$ increases by $1/2$.
+As shown in Table 3, if the goal is to maximize the probability of guessing $s$, based on one query to either the unencoded or encoded oracles, then for $n\geq 11$ the unencoded oracle is better.
+On the other hand, if the goal is to minimize the expected Hamming distance to $s$, then it is better to use the encoded oracle.
+These statements can both hold because the failure modes of the encoded and unencoded algorithms are different (Fig.
+9).
+Intuitively, an error on the ancilla qubit in the unencoded algorithm typically results in a string like $1^{n/2}0^{n/2}$, contributing $n/2$ to the distance from $s$.
+However, with the encoded algorithm a faulty ancilla block more likely leads to $1^{n/2}\star^{n/2}$, where $\star$ denotes an undecodeable block; this contributes only $n/4$ to the expected distance.
+Typical errors in the encoded algorithm are heralded, so they contribute less to the expected distance.
+Either quantum oracle is much better than the classical oracle.
+One classical query can only return one bit of $s$, so the probability of guessing $s$ is $\Pr[\text{guess}]=1/2^{n-1}$ and the expected distance to $s$ is $\operatorname{\operatorname{E}}[\text{Hamming}]=\tfrac{n-1}{2}$.
+Table 3: Performance of unencoded and encoded Bernstein-Vazirani algorithms with one oracle query.
+$\Pr[\text{guess}]$ is the probability of guessing the secret $s$, and $\operatorname{\operatorname{E}}[\text{Hamming}]$ the expected distance from $s$.
+For a classical oracle, $\Pr[\text{guess}]=1/2^{n-1}$ and $\operatorname{\operatorname{E}}[\text{Hamming}]=\tfrac{n-1}{2}$.
+  Unencoded baseline Encoded algorithm
+$n$ $\Pr[\text{guess}]$ $\operatorname{\operatorname{E}}[\text{Hamming}]$ $\Pr[\text{guess}]$ $\operatorname{\operatorname{E}}[\text{Hamming}]$
+7 67(4)0.94 75(2)0.48
+11 69(4)1.04 66(3)0.90
+15 61(4)1.58 51(3)1.45
+19 49(3)2.44 38(2)2.29
+23 35(2)3.42 30(2)2.97
+27 28.2(7)4.12 18.5(5)4.05
+[…]
+Outlook
+In this work, we have created large entangled logical states and implemented an algorithm fault tolerantly on encoded qubits, with better-than-physical performance.
+These results highlight the features and promise of neutral atoms as a platform for reliable fault-tolerant quantum computing.
+They are enabled by the combination of scalability and high-fidelity operations inherent to neutral atoms, with full qubit connectivity.
+All of the experiments highlight the usefulness of loss conversion and imaging.
+The encoded experiments demonstrate that logical circuits can correct for and operate effectively with atom loss, which will be a key feature of large-scale neutral atom quantum computers.
+This work, and other recent work with neutral-atom qubits BEG+ (24); RRJ+ (24), superconducting qubits GC (24) and trapped ion qubits PSR+ (24); RAC+ (24), highlights a critical advance in the field of quantum computation: the transition from noisy, intermediate-scale quantum computations with physical qubits to computational advances achieved with logical qubits.
+After a relatively short time from its inception, the neutral atom quantum computing platform has shown significant advancement.
+We expect further advances in 2Q gate fidelities and significant scale up towards 10 0001000010\,00010 000 qubits, including mid-circuit measurement NC+ (23) and continuous atom reloading NKC+ (24).
+Larger scales will unlock nonlocal error-correcting codes with high encoding rates XZZ+ (24).
+The combination of a hardware-optimized qubit virtualization system, complete with efficient error correction, with a programmable neutral atom quantum processor is a promising path to large-depth, logical computation for scientific quantum advantage.
+Acknowledgments
+We thank Zulfi Alam, David Bohn, Jeongwan Haah, Chetan Nayak, Pasi Kostamo, Vadym Kliuchnikov, Dennis Tom, Matthias Troyer, and Jason Zander for their support and feedback during this project.
+Appendix A Methods
+A.1 Register loading and atom reuse
+We load atoms into the register using the methods described in NKC+ (24), providing more than enough qubits for the circuit (typically filling 320 register sites to approximately 99% filling fraction over up to 14 loading cycles).
+The register sites are configured in eight pairs of columns, matching the shape of the IZ.
+This provides “highways” through which atoms can be moved with less risk of disturbing other atoms in the register.
+By loading more register sites than are used in a given circuit, we can improve data rates by reusing atoms in multiple circuit instances.
+To do this, we determine which atoms were lost in a circuit, and refill those sites using extra atoms, then perform the rest of our state preparation protocol (described in the following section).
+This can be repeated as long as atoms remain in the register.
+A.2 State preparation and measurement
+Our platform benefits from the ability to determine successful loading of our desired qubit array before circuit execution, and to separately determine the survival and final state of an atom after the circuit.
+We achieve this by performing three images per circuit instance, in addition to images used to initially arrange atoms into the qubit register.
+A first “preselect” image determines whether the atoms were successfully loaded into the register before executing the circuit.
+A second “readout image” comprises the state readout of the circuit.
+A third “survival” image determines whether an atom is still present after the circuit.
+In all images, we measure the presence of an atom in state ${|1\rangle}$.
+The pre-select and survival images are combined with optical pumping into ${|1\rangle}$, so only determine the site occupancy.
+In the readout image, the optical pumping is omitted and state selectivity is maintained.
+Following the preselect image, we perform gray-molasses cooling and optically pump into state ${|0\rangle}$.
+Aspects of our imaging procedure have been described in NC+ (23); NKC+ (24).
+The imaging is performed in a cavity-enhanced 784 nm wavelength 3D lattice by collecting the photons scattered from the cycling transition from ${|1\rangle}$ to ${}^{3}P_{1}{|3/2,3/2\rangle}$.
+Optical pumping and gray-molasses cooling are also performed in the cavity-enhanced lattice.
+Between imaging and circuit execution, atoms are transferred between the lattice and register tweezers.
+Typical classification errors, loss probabilities, and spin-flip probabilities for the imaging are provided in NKC+ (24).
+Here, we also characterize state preparation and measurement errors in a manner relevant to our circuit execution by simply preparing either state ${|0\rangle}$ or ${|1\rangle}$, and then measuring the state of the atoms, conditioned on the presence of an atom in the survival image.
+We obtain combined preparation and readout errors for ${|0\rangle}$ and ${|1\rangle}$ of $1.1(1)\times 10^{-3}$ and $1.6(1)\times 10^{-3}$, respectively.
+Histograms of the recorded photon counts provide insight into the nature of the errors (Fig.
+10).
+For both prepared states, we see clear evidence of a two-part distribution, indicating that the dominant contribution to the error is that the atom actually ends up in the wrong state, rather than overlap of the photon count distribution of dark and bright atoms.
+This is likely due to a combination of imperfect optical pumping and spin flips that occur within the image when measuring ${|1\rangle}$.
+(Spin flips are primarily due to population of ${}^{3}P_{1}$, so are suppressed when measuring the dark state ${|0\rangle}$.)
+Figure 10: a, Timing diagram of imaging and optical pumping used for circuit execution.
+Optical pumping is used both to provide state insensitivity for preselect and survival images, and for state preparation.
+b, Photon count histograms recorded after preparing ${|0\rangle}$ (left) and ${|1\rangle}$ (right).
+The dashed vertical line indicates the count threshold used to distinguish the two states.
+Error rates for ${|0\rangle}$ and ${|1\rangle}$ are 0.0011(1) and 0.0016(1), respectively.
+A.3 Atom movement
+Arbitrary gate connectivity in our system is obtained by moving selected atoms from the register into an interaction zone (IZ) where two-qubit gates are performed (see Fig.
+1).
+Movement is performed using a set of tweezers generated by a pair of crossed AODs.
+Atoms within a single row can be moved in parallel, provided that the move profile does not involve parasitic tweezer spots generated by intermodulation tones crossing the position of the intentional tweezers used for movement.
+In practice, we guarantee this condition by requiring parallel moves of more than two atoms to be shape-preserving (stretching and translation of the tweezer locations only).
+Moves are broken up into straight segments, with the movement velocity along each segment following a sinusoidal profile, with a total duration scaled to maintain a fixed peak acceleration (as the square root of the length of the segment).
+Move durations are $0.12\sqrt{L}$ ms, where $L$ is the length of the move in units of the 3.3 $\mu$m grid spacing of the register.
+Atoms are handed between static and rearrangement tweezers by ramping the depth of the rearrangement tweezers over 0.4 ms.
+For up to 50 moves, we observe a typical loss probability of $5(1)\times 10^{-4}$.
+For more moves, loss per move increases.
+By performing sideband spectroscopy on the 1S0 to 3P0 transition after a variable number of moves, we find a heating rate of 0.02(5) quanta along the propagation direction of our clock addressing laser.
+This heating rate is too low to explain the increased rate of atom loss above 50 moves, making axial heating a more likely culprit for the excess loss.
+In our experiments, atom movement is scheduled using a combination of manual and algorithmic techniques, including the use of graph and annealing algorithms to minimize the number of atom movement operations.
+A.4 Single-qubit gate performance
+Our single-qubit (1Q) gate scheme is described in detail in Ref.
+MSS+ (24), and in that work characterized within a small array.
 These gates are performed using two-photon Raman transitions driven by two independently controlled and site-resolved laser beams detuned roughly 10 GHz from the 3P1, $F=1/2$ manifold.
+Single-qubit operations are performed in parallel on rows of atoms within the register, with the pulse area, phase, and presence of the gates controlled individually on each site.
 We calibrate an $S_{X}$ gate on each site of the register, and perform virtual $Z$ rotations of arbitrary angle by changing the relative phase of the two drive lasers on subsequent gates.
-We characterize the performance of the single-qubit operations using a Clifford randomized benchmark sequence.
-At depths of up to 500 Clifford gates, we find an average fidelity of 99.85(2)per Clifford gate, averaged across three rows of the register spanning those used in this work.
-Because the nuclear spin qubit is well-isolated from differential light-shifts and has low sensitivity to magnetic fields, we do not automatically apply dynamical decoupling sequences for the circuits presented in this work.
-Two-qubit controlled-phase (CZ) gates are performed in the IZ on up to eight pairs of atoms in parallel.
+Figure 11: Characterization of single-qubit gates.
+a, Array-averaged results of Clifford randomized benchmarking, indicating a per-Clifford-gate error of $6.7\times 10^{-4}$.
+b, Characterization of the spread in per-Clifford-gate errors on different sites within six rows sampled within the register.
+Here, we characterize the 1Q gates in the larger register array used in this work.
+As in MSS+ (24), we perform Clifford randomized benchmarking.
+We consider qubits occupying six rows that span those used in our circuits and extract an average error rate per Clifford gate of $6.7\times 10^{-4}$ (see Fig.
+11a).
+The spread in gate fidelity for different qubit sites is illustrated in Fig.
+11b.
+In addition to errors occurring during gates, idle errors can also be relevant.
+Using Ramsey spectroscopy, we measure a root-mean-squared spread in the qubit frequencies of 0.52 Hz after software compensation of shifts due to magnetic field gradients and average light shifts.
+The ensemble-averaged coherence exhibits a Gaussian $1/e$ decay time of 3.5(17) seconds.
+Uncorrected shifts are likely dominated by inhomogeneous light shifts from the trapping tweezers, which we measure to be typically on the Hz scale for the register tweezers.
+Additionally, the rearrangement tweezers lead to shifts on the several-Hz scale, which we calibrate and cancel in software.
+These shifts are not fully understood, but may result from finite helicity in the tweezer polarization.
+These frequency shifts are small enough that we do not apply dynamical decoupling sequences for the circuits in this work.
+A.5 Two-qubit gate implementation and performance
+Two-qubit controlled-phase (CZ) gates are performed in the IZ on up to eight pairs of atoms in parallel, using the same scheme described in MSS+ (24).
 These gates are mediated by excitation to a highly-excited Rydberg state ($65^{3}\mathrm{S}_{1}$).
@@
 Polarization selectivity, along with the narrow linewidth of the transition and slow Rabi rate (10 kHz) relative to the qubit frequency (376 kHz) ensure that only one qubit state is excited to the metastable state.
-From the metastable state, atoms are excited to the Rydberg state using a UV laser pulse (total duration 110 ns) that implements a gate similar to the time-optimal controlled-phase (CZ) gate described in JP (22).
+From the metastable state, atoms are excited to the Rydberg state using a UV laser pulse (total duration 110 ns) that implements a gate similar to the time-optimal controlled-phase (CZ) gate described in JP (22); PWJ+ (22); EBK+ (23).
 While the IZ traps are equally confining for the ground and metastable states, the register traps are near a tune-out wavelength for the metastable states, providing near-zero trapping potential.
-Thus, any gate errors that lead to population being left in the metastable state (such as some coherent population errors on either stage of the two-photon transition) are converted to loss and can be detected.
+Thus, any gate errors that lead to population being left in the metastable state (such as some coherent population errors on either stage of the two-photon transition) are converted to loss and can be detected (see Sec.
+A.6).
 This “erasure conversion” proves useful in identifying the location of errors WKPT (22).
-We quantify the performance of our two-qubit gates using several variations on randomized benchmarking.
-In all presented cases, data is taken in parallel over the eight pairs of sites in the IZ, and averaged.
-To isolate the performance of the two-qubit gates themselves, in the absence of any errors that may be incurred through atomic movement, we perform interleaved randomized benchmarking (IRB) on static atoms within the IZ.
+We quantify the performance of two-qubit gates using several variations on randomized benchmarking.
+To isolate the performance of the gates themselves, without errors from atom movement, we perform interleaved randomized benchmarking (IRB) on static atoms within the IZ.
 In this protocol, we compare the return probability after a given number of random two-qubit Clifford operations (plus a return pulse calculated to return to a desired measurement state) to the same circuit with an additional CZ gate interleaved after each Clifford gate (and with an appropriately modified return pulse).
 By comparing the probability of returning to the measurement state versus the depth of the Clifford circuit with and without the added CZ gates, we isolate the error associated with the CZ gate alone, averaged over the two-qubit Hilbert space.
-Here, the atoms are kept in the IZ for all operations (unlike typical circuits, single qubit gates are applied in the IZ for this protocol).
+Here, the atoms are kept in the IZ for all operations (including 1Q gates, unlike for typical circuits).
 Fits to IRB data taken out to depth 34 indicate an infidelity per CZ gate of 0.56(15)and additional loss and population leakage outside of the qubit subspace of 0.27(6)%.
-Table 2: Randomized benchmark infidelities, post-selected on an atom (1Q) or pair of atoms (2Q) remaining in the qubit subspace, and probability of one or both atoms being lost from the qubit subspace.
-Benchmarking sequences are described in the main text.
-  Benchmark protocol Infidelity Loss + Leakage
-1Q Clifford RB 0.15(2)0.007(1)%
-2Q Static IRB 0.56(15)0.27(6)%
-2Q Echoed RB with moves 0.4(1)0.5(1)%
-2Q Echoed RB without moves 0.35(5)0.24(6)%
-In actual circuits, atoms are moved between the register and IZ between single-qubit operations and two-qubit operations, to allow arbitrary connectivity.
+In actual circuits, atoms are moved between the register and IZ between 1Q operations and 2Q operations.
 Because atom movement can cause heating, which may impact the performance of all following gates, IRB does not cleanly isolate the performance of the CZ gate when atoms are moved between arrays.
-For this, we rely on a variant of randomized benchmarking described in MSS+ (24) and similar to that in EBK+ (23), where concatenated circuit blocks each consist of a random single-qubit operation and two CZ gates separated by an echo pulse.
-A final gate sequence returns atoms to the measurement basis.
-We isolate the contribution from the two-qubit gates (along with any loss or infidelity due to associated movements, up to single-qubit phases which are canceled by the echo pulses) by subtracting the per-depth infidelity associated with the single-qubit operations.
-From this, we extract a per-CZ gate error of 0.4(1)and loss of 0.5(1)%.
+For this, we rely on a variant of randomized benchmarking described in MSS+ (24) and similar to that in EBK+ (23), where concatenated circuit blocks each consist of a random 1Q operation and two CZ gates separated by an echo pulse.
+A final 1Q gate returns atoms to the measurement basis.
+We isolate the contribution from the 2Q gates (along with any loss or infidelity due to associated movements, up to single-qubit phases which are canceled by the echo pulses and characterized separately) by subtracting the per-depth infidelity associated with the 1Q operations.
+From this, we extract a per-CZ gate error of 0.39(10)and loss of 0.75(10)%.
+We perform this characterization with atoms originating in different rows within the register, as shown in Fig.
+12, and do not observe a significant dependence on origin row or column.
 For comparison, the same procedure applied to atoms that remain static in the IZ for all operations yields a per-gate error rate of 0.35(5)and loss of 0.24(6)%.
-Table 2 summarizes the results of these benchmarks.
-The performance of our gates is sensitive to relative alignment of gate lasers and optical potentials, and other factors that may drift on the timescale of hours to days.
-The values presented here represent typical values over the time period in which the circuits presented in this work were run.
-In our experiments, atom movement is scheduled using a combination of manual and algorithmic techniques, including the use of graph and annealing algorithms to minimize the number of atom movement operations.
-III Four-qubit codes
+Figure 12: Characterization of CZ gate performance on moved atoms.
+a, Echoed randomized benchmarking circuit applied with CZ gates omitted.
+Orange, green and blue points represent survival, fidelity for atoms that survive, and combined survival and fidelity, respectively.
+Lines represent exponential fits used to extract per-block error and loss rates.
+b, The same quantities with CZ gates added in (2 CZ gates per block).
+c, Measured infidelity (green) and loss (orange) for different qubit origin locations within the register.
+Statistical fluctuations may yield negative infidelity estimates because per-block error rates with CZ gates omitted are subtracted from those with CZ gates included.
+A.6 Conversion of leakage to loss
+Gate errors in the device, coherent or incoherent, can lead to population and phase errors within the qubit subspace, as well as leakage into other states.
+We operate the machine in a mode that converts population leakage from the qubit subspace into loss, which can then be identified independently from qubit state.
+This is particularly relevant in the context of 2Q gates, where population errors on the 1S0 to 3P0 or 3P0 to Rydberg state transitions can lead to population leakage.
+Conversion of these errors into loss is achieved by operating our register tweezers at a wavelength near 423 nm where 3P0 is nearly untrapped.
+(We could choose to depump the 3P0 states after each 2Q gate to avoid this loss, but find the conversion of leakage into a detectable error highly desirable.)
+We characterize the effectiveness of the conversion from leakage to loss for 3P0 and the Rydberg state by preparing atoms in the register, moving them to the IZ, exciting them to the relevant state, moving them back to the register and then measuring the state and presence of the atoms.
+The metastable 3P2 states are anti-trapped in our IZ and rearrangement tweezers, and we do not repump these states.
+We observe greater than 99% of atoms to be lost when exciting to the clock state, and roughly 70% of atoms to be lost when exciting to the Rydberg state.
+Appendix B Four-qubit codes
 The $\llbracket 4,2,2\rrbracket$ code VGW (96); GBP (97) is a stabilizer code Got (98) with stabilizer generators
@@
 Observe that encoded ${|00\rangle}$ is the four-qubit cat state $\tfrac{1}{\sqrt{2}}({|0000\rangle}+{|1111\rangle})$.
-As shown in Figs.
-4 and 5, an extra ancilla qubit is needed to prepare this state fault tolerantly Sho (96).
+As shown in Fig.
+5, an extra qubit is needed to prepare this state fault tolerantly Sho (96).
 The $\llbracket 4,1,2\rrbracket$ code is actually the same code, but only using one of the two logical qubits.
@@
 Finally, using $H=S_{Z}S_{X}S_{Z}$, applying $S_{X}\otimes S_{X}^{\dagger}\otimes S_{X}^{\dagger}\otimes S_{X}$ implements a logical dual CZ gate, $(H\otimes H)\,\text{CZ}\,(H\otimes H)$.
-IV Encoded $24$-qubit cat state
-An $n$-qubit cat state can be written as $\tfrac{1}{\sqrt{2}}({|0^{n}\rangle}+{|1^{n}\rangle})$.
-Here, we prepare an encoded cat state on $24$ logical qubits, encoded with the $\llbracket 4,2,2\rrbracket$ code into $48$ physical qubits.
-The circuit we implement is explained in Fig.
-4.
-Starting with a logical circuit to prepare $\tfrac{1}{\sqrt{2}}({|0^{24}\rangle}+{|1^{24}\rangle})$, we fault-tolerantly encode it into the $\llbracket 4,2,2\rrbracket$ code.
-We then arrange the qubits, four code blocks in each row with the ancilla qubits at the end, and schedule the gates and qubit movements as in Fig.
-5.
-The unencoded baseline experiment prepares a $24$-qubit cat state using a separately optimized circuit, in Fig.
-6.
-[…]
-Figure 5: A portion of the encoded circuit from Fig.
-4 prepares an encoded four-qubit cat state.
-The gates in this circuit need to be scheduled.
-Two-qubit CZ gates are applied in the interaction zone, up to eight gates in parallel, and all the one-qubit gates on a particular qubit are applied at its initial location in the storage zone.
-One-qubit gates on a subset of qubits in the same row can be applied in parallel.
-Two atoms can be moved in parallel, but their paths cannot cross.
-Note that in this example, for clarity the two code blocks are arranged adjacent to each other in the storage zone, whereas in Fig.
-4 they are interleaved.
-Figure 6: Circuit to prepare a $24$-qubit cat state, unencoded. One-qubit gates are applied at an atom’s home register location.
-The cat state has genuine multipartite entanglement, meaning that it is entangled across every bipartition of the $n$ qubits.
-Any state that has fidelity greater than $1/2$ with the ideal cat state also has genuine multipartite entanglement GS (10).
-However, states with $\leq 1/2$ fidelity may not even be entangled; the classical mixture $\tfrac{1}{2}({{|0^{n}\rangle}\!{\langle 0^{n}|}}+{{|1^{n}\rangle}\!{\langle 1^{n}|}})$ has fidelity $1/2$ with the ideal cat state, but no entanglement.
-Our experiments, as well as HDHL (24); RAC+ (24), measure $X$ basis and $Z$ basis error rates, $p_{X}$ and $p_{Z}$.
-The fidelity to the ideal cat state is between $1-p_{X}-p_{Z}$ and $1-\max(p_{X},p_{Z})$, essentially because possible $Y$ errors contribute to both $p_{X}$ and $p_{Z}$; this is the reason for the fidelity ranges in Table 1.
-We find that for the encoded cat state, using error detection and loss correction, $p_{X}+p_{Z}=26.6^{+3.3}_{-3.1}\%$, while for the unencoded baseline $p_{X}+p_{Z}=42.0^{+11.3}_{-10.1}\%$.
-The results are summarized in Fig.
-1 and Table 3.
-While the encoding is clearly beneficial, interestingly, it helps reduce $p_{X}$ much more than $p_{Z}$.
-The baseline error bar is large because rather than averaging over all trials, we chose for a challenging baseline the best of $27$ separate runs, each consisting of 300300300300400400400400 unencoded trials with $X$ and $Z$ basis measurement; Fig.
-7 gives more details.
-Figure 8 shows how conditioning on at most $k$ lost qubits, for various values of $k$, allows a smoother interpolation, in both acceptance probability and total error probability, between the different results in Table 3.
-Finally, Fig.
-9 shows the different machine runs that combine to give the data in Table 3, plotted conditioned on no initial atom loss.
-The spread and correlations in the figure suggest some machine parameter drift.
-Figure 7: Each trial of the unencoded baseline $24$-qubit cat state experiment included two cat states, one measured in the $X$ basis and the other measured in the $Z$ basis.
-Shown here are the baseline runs, each consisting of 300300300300400400400400 trials.
-Only cat states with no qubit loss are included, on average $135.3$ and $161.0$ trials for $X$ and $Z$ basis measurements, respectively.
-The high variance in the observed error rates complicates reporting a single baseline $X$ plus $Z$ basis total error rate.
-To be conservative, disadvantaging the encoded cases, we decided to report the best single run as a baseline.
-The red data point has the least total error.
-In the $X$ basis, $26$ out of $100$ cat states had a logical error, and in the $Z$ basis, $19$ out of $121$ cat states had a logical error.
-The baseline error bars in Fig.
-1 are large because they are based on only this run, rather than on the average of all the runs in this plot.
-Figure 8: Data from the encoded $24$-qubit cat state experiment, selecting only trials within specified loss criteria.
-In general, attempting to correct for more loss leads to a higher acceptance probability, but also a higher error rate.
-In each trial, an atom may be missing before the circuit even begins, a loading loss, or it may be lost during the circuit.
-The red points show the total $X$ basis plus $Z$ basis error rate, plotted against the average of the $X$ and $Z$ acceptance probabilities, based on either all 29 7602976029\,76029 760 trials, only those 17 0081700817\,00817 008 trials with no initial loss, or only those $858$ trials with no loss at all.
-For the blue points, we select only trials in which the total number of lost qubits, including load loss and circuit loss, is at most a certain threshold.
-Note that Fig.
-9 only includes trials with no initial loss, and the acceptance probabilities in Fig.
-9 are conditioned on no initial loss, so the normalization is different.
-(Figure 7 for the baseline only includes trials with no loss at all.)
-Table 3: Data for the $24$-qubit cat state experiments.
-An experiment is accepted if it is neither prerejected nor postrejected.
-For the encoded experiments, a trial is “prerejected” if it is rejected during the initial state preparation because one of the ${|\overline{00}\rangle}=\tfrac{1}{\sqrt{2}}({|0^{4}\rangle}+{|1^{4}\rangle})$ states fails verification.
-A surviving trial is “postrejected” if a code block cannot be decoded, either because of a detected error or because of two or more lost qubits in the block.
-We have also filtered out the encoded data with no initial losses, and the data with no losses at all.
-For the unencoded baseline, a trial is prerejected if there are any initial losses and a surviving trial is postrejected if there are any circuit losses.
-As explained in Fig.
-7, we have selected the best of the unencoded runs, in order to create a challenging baseline.
-  Experiment Basis Trials Prerejected Postrejected $\Pr[\text{accept}]$ Errors Error rate
-Cat-$24$ encoded: all data $X$ 14 4001440014\,40014 400 10 1591015910\,15910 159 $3026$ $8.4^{+0.5}_{-0.4}\%$ 148 $12(2)\%$
-$Z$ 15 3601536015\,36015 360 11 0901109011\,09011 090 3283 $6.4(4)\%$ 170 $17(2)\%$
-no initial loss $X$ $8433$ 5371 2122 $11.1(7)\%$ 115 $12(2)\%$
-$Z$ $8575$ 5589 2185 $9.3(6)\%$ 115 $14^{+3}_{-2}\%$
-no loss $X$ 395 66 221 $27^{+5}_{-4}\%$ 9 $8^{+6}_{-4}\%$
-$Z$ 463 98 220 $31(4)\%$ 2 $1^{+3}_{-1}\%$
-Cat-$24$ unencoded baseline $X$ 300 87 113 33(5)26 $26^{+8}_{-9}\%$
-$Z$ 300 82 97 $40^{+6}_{-5}\%$ 19 $16^{+7}_{-6}\%$
-(a) $X$ basis runs
-(b) $Z$ basis runs
-Figure 9: Encoded $24$-qubit cat state data from 45 $X$ basis measurement runs, and 48 $Z$ basis measurement runs.
-This data was taken overnight, while the machine ran unattended.
-Each blue point gives the acceptance probability, conditioned on no initial loss, and logical error rate for a run of, on average, $187$ trials for the $X$ basis data, and $179$ trials for the $Z$ basis data.
-(We attempted $320$ trials, but some trials failed to load all atoms.) For clarity, only some error bars are shown.
-The black squares are the averages over all the data, the numbers we report.
-We interpret the negative correlation between acceptance probability and logical error rate as evidence of some machine parameter drift between runs, despite automatic recalibrations.
-V Encoded Bernstein-Vazirani algorithm
-Consider a black box oracle $\{0,1\}^{n}\rightarrow\{0,1\}$, with a secret string $s\in\{0,1\}^{n}$, that on query $x$ returns $s\cdot x=s_{1}x_{1}\oplus\cdots\oplus s_{n}x_{n}$, with $\oplus$ denoting addition modulo $2$.
-Classically, $n$ queries to the oracle are required and sufficient to learn $s$.
-The Bernstein-Vazirani algorithm BV (97); CEMM (98) makes a single query to the quantum version of this oracle, $U_{s}=\sum_{x\in\{0,1\}^{n},b\in\{0,1\}}{{|b\oplus(s\cdot x)\rangle}\!{\langle b|}}\otimes{{|x\rangle}\!{\langle x|}}$, to learn $s$ through the following steps:
-${|-,+^{n}\rangle}$ $\overset{U_{s}}{\longmapsto}\tfrac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}{(-1)^{s\cdot x}{|-,x\rangle}}$ $\overset{I\otimes H^{\otimes n}}{\longmapsto}{|-,s\rangle}\enspace.$
-This classic algorithm was designed to prove an oracle separation between the classical randomized and quantum polynomial-time complexity classes $\mathsf{BPP}$ and $\mathsf{BQP}$.
-[…]
-Figure 10: The $n=19$ Bernstein-Vazirani algorithm involves $n+1=20$ logical qubits.
-In our experiment, we test the most challenging case, in which the secret string is $1^{n}$, so there are $n$ logical CNOT gates.
-Encoded into the $\llbracket 4,1,2\rrbracket$ code, each logical qubit becomes a block of four atoms.
-There are $4(n+1)$ atoms and $26n+$ 2 physical CZ gates.
-Four code blocks are arranged in each of the storage zone’s top five rows.
-The gates are scheduled so that the first block, once initialized, stays in the interaction zone.
-As transversal CZ gates are applied into block $j$, block $j+1$ is initialized.
-The unencoded algorithm accepts if no measured atom is lost, and is successful if the measured string is $1^{n}$.
-The encoded algorithm tolerates losses, but rejects if any of the last $n$ code blocks is not decodable.
-Figure 2 summarizes the results, with more details in Table 4.
-The encoded algorithms have a statistically significantly higher success probability than the unencoded baseline, albeit with a lower acceptance rate.
-Table 4: Data for the Bernstein-Vazirani experiments.
-The encoded experiment has no preselection.
-A surviving trial is rejected if one of the last $n$ code blocks cannot be decoded, either because of a detected error or from too many lost qubits.
-(The first code block is ignored.) For the unencoded baselines, a trial is rejected if a measured qubit is lost.
-      All data
-$n$ Version Trials Rejected $\Pr[\text{accept}]$ Errors Error rate
-7 encoded 1200 409 $66(3)\%$ 18 $2.3^{+1.2}_{-0.9}\%$
-baseline 600 224 $63(4)\%$ 57 $15^{+4}_{-3}\%$
-11 encoded 1200 547 $54(3)\%$ 22 $3^{+2}_{-1}\%$
-baseline 600 196 $67(4)\%$ 65 $16^{+4}_{-3}\%$
-15 encoded 1200 740 $38(3)\%$ 40 $9^{+3}_{-2}\%$
-baseline 600 243 $59(4)\%$ 73 $20(4)\%$
-19 encoded 1800 1359 $25(2)\%$ 43 $11(3)\%$
-baseline 800 457 $43(3)\%$ 90 $26^{+5}_{-4}\%$
-23 encoded 2400 1993 $17^{+2}_{-1}\%$ 67 $16^{+4}_{-3}\%$
-baseline 3160 2379 $25^{+2}_{-1}\%$ 212 $27(3)\%$
-27 encoded 26 0402604026\,04026 040 23 7892378923\,78923 789 $8.6(3)\%$ 498 $(2)\%2$
-baseline 15 6001560015\,60015 600 12 8771287712\,87712 877 $17.5(6)\%$ 857 $31(2)\%$
-Trials with no initial loss
-$n$ Version Trials Rejected $\Pr[\text{accept}]$ Errors Error rate
-7 encoded 732 233 $68(3)\%$ 7 $1.4_{-0.8}^{+1.3}\%$
-baseline 425 76 $82_{-4}^{+3}\%$ 30 $9(3)\%$
-11 encoded 471 181 $62(4)\%$ 4 $1.4_{-1.0}^{+1.8}\%$
-baseline 540 142 $74(4)\%$ 59 $15_{-3}^{+4}\%$
-15 encoded 543 308 $43(4)\%$ 15 $6_{-3}^{+4}\%$
-baseline 526 176 $67(4)\%$ 66 $19(4)\%$
-19 encoded 537 369 $31(4)\%$ 16 $10_{-4}^{+5}\%$
-baseline 623 289 $54(4)\%$ 81 $24_{-4}^{+5}\%$
-23 encoded 800 647 $19(3)\%$ 25 $16_{-5}^{+6}\%$
-baseline 2173 1409 $35(2)\%$ 195 $26(3)\%$
-27 encoded 4981 4443 $10.8_{-0.8}^{+0.9}\%$ 104 $19(3)\%$
-baseline 10 3581035810\,35810 358 7695 $25.7(8)\%$ 797 $30(2)\%$
-Computing the success probability conditioned on acceptance makes sense if we treat the Bernstein-Vazirani algorithm as a generic quantum algorithm, as done above.
-However, it was originally proposed as a quantum query algorithm, in which minimizing the number of oracle queries was of key importance.
-In this light, we can look at the data and ask whether with only one oracle query allowed (one trial), the encoded circuit still beats the unencoded circuit.
-In short, the answer is probably not.
-In this case, neither the encoded nor unencoded algorithms can condition on acceptance, and with only one trial it requires guessing the secret string.
-In particular, with a lost unencoded qubit or an undecodable code block the probability of guessing the corresponding bit is only $1/2$, and the expected Hamming distance of the guess away from the secret string increases by $1/2$.
-As shown in Table 5, if you are only allowed to make one query to either the unencoded or encoded oracles, and if your goal is to maximize the probability of guessing the secret string $s$, then for $n\geq 11$ it is better to use the unencoded oracle.
-On the other hand, if your goal is to minimize the expected Hamming distance of your guess from $s$, then it is better in all cases to use the encoded oracle.
-These statements are not contradictory.
-They hold because the failure modes of the encoded and unencoded algorithms are quite different, as shown in Fig.
-11.
-Intuitively, an error on the ancilla qubit in the unencoded algorithm typically results in a string like $1^{n/2}0^{n/2}$ and contributes on average $n/2$ to the expected Hamming distance from $s$.
-On the other hand, an error on the ancilla code block in the encoded algorithm often results in a string like $1^{n/2}\star^{n/2}$, where $\star$ denotes an undecodeable code block where one must guess a bit of $s$; this contributes only an average of $n/4$ to the expected Hamming distance.
-Typical errors in the encoded algorithm are heralded, so they contribute less to the expected Hamming distance error metric.
-One query to either quantum oracle is much better than one query to the classical oracle.
-One classical query can only return one bit of $s$, so the probability of guessing $s$ is $\Pr[\text{guess}]=1/2^{n-1}$ and the expected Hamming distance to $s$ is $\operatorname{\operatorname{E}}[\text{Hamming}]=\tfrac{n-1}{2}$.
-Table 5: Performance of unencoded and encoded Bernstein-Vazirani algorithms with a single oracle query.
-$\Pr[\text{guess}]$ is the probability of guessing the secret string based on the measured outputs, and $\operatorname{\operatorname{E}}[\text{Hamming}]$ the expected distance of the guess away from the secret string.
-This analysis is based on all trials, including ones with initial loss.
-For a classical oracle, $\Pr[\text{guess}]=1/2^{n-1}$ and $\operatorname{\operatorname{E}}[\text{Hamming}]=\tfrac{n-1}{2}$.
-  Unencoded baseline Encoded algorithm
-$n$ $\Pr[\text{guess}]$ $\operatorname{\operatorname{E}}[\text{Hamming}]$ $\Pr[\text{guess}]$ $\operatorname{\operatorname{E}}[\text{Hamming}]$
-7 67(4)0.94 75(2)0.48
-11 69(4)1.04 66(3)0.90
-15 61(4)1.58 51(3)1.45
-19 49(3)2.44 38(2)2.29
-23 35(2)3.42 30(2)2.97
-27 28.2(7)4.12 18.5(5)4.05
-[…]
-VI Repeated loss correction and transversal operations for the $\llbracket 4,2,2\rrbracket$ code
-VI.1 Repeated loss correction with CZ and dual CZ logical gates
-The circuits considered so far have only used loss correction and error detection at the very end, when decoding the final measurements.
-However, it’s important to be able to detect and correct errors between computational steps of an algorithm, to enable deeper computation.
-To that end, we explore using loss correction in the middle of a larger circuit, demonstrating repeated fault-tolerant loss correction.
+Appendix C Repeated loss correction and transversal operations for the $\llbracket 4,2,2\rrbracket$ code
+C.1 Repeated loss correction with CZ logical gates
+The circuits considered so far have only used loss correction and error detection at the very end, when interpreting the final measurements.
+However, it is important to be able to detect and correct errors between the steps of an algorithm, to enable deeper computation.
+To that end, we demonstrate repeated fault-tolerant loss correction.
 In order to make the circuits more interesting, between rounds of loss correction we apply some of the fault-tolerant logical operations given in Sec.
-III.
+B.
+Repeated error detection with distance-two codes has been demonstrated on ion traps, including three rounds on eight logical qubits SBA (24), and 15 rounds on four logical qubits YDKR (24).
+Demonstrations based on Gottesman’s protocol Got (16) for fault tolerance with the $\llbracket 4,2,2\rrbracket$ code, but without including repeated error detection, include WWJ+ (18); HF (19); Vui (18); LGL+ (17); CCNH (21).
+Repeated error correction, with higher-distance codes, has also been demonstrated on on trapped-ion processors RBL+ (21); PBP+ (24); PSR+ (24); RAC+ (24) and superconducting processors KLR+ (22); Goo (23); GC (24); PNH+ (24).
 Figure 13 gives a circuit for fault-tolerant error detection for the $\llbracket 4,2,2\rrbracket$ code.
@@
 Furthermore, the circuit in Fig.
-12(b) replaces two of the data qubits with fresh qubits, so the full data block can be refreshed every two rounds of error detection.
+13(b) replaces two of the data qubits with fresh qubits, so the full data block can be refreshed every two rounds of error detection.
 This allows us to use the $\llbracket 4,2,2\rrbracket$ code to correct a lost qubit during a computation, not just at the very end as in our logical cat state procedure.
-Figure 12: The top circuit (a) measures the parity of four qubits, i.e., the syndrome of $ZZZZ$.
-However, it is not fault tolerant; a single $Z$ fault on the ancilla can spread undetected to a weight-two $Z$ error.
-The next circuit (b) remedies this, and simultaneously measures $XXXX$, using a trick from Rei (20).
+Figure 13: a, Circuit to measure the parity of four qubits, i.e., the syndrome of $ZZZZ$.
+It is not fault tolerant; a single $Z$ fault on the ancilla can spread undetected to a weight-two $Z$ error.
+b, The next circuit remedies this, and simultaneously measures $XXXX$, using a trick from Rei (20).
 Now the bad $Z$ fault will be detected, or “flagged,” by the $XXXX$ measurement, and similarly the $ZZZZ$ measurement will flag a bad $X$ fault.
-This circuit gives an efficient, fault-tolerant error detection procedure for the $\llbracket 4,2,2\rrbracket$ code.
+This gives an efficient, fault-tolerant error detection procedure for the $\llbracket 4,2,2\rrbracket$ code.
 However, it does not correct for a lost qubit.
-The bottom circuit (c) fixes this, halfway, by switching the direction of the final two CNOT gates.
-(This is equivalent to inserting swap-based “leakage-detection units” CBO+ (24).) The last two data qubits are replaced with fresh ancillas, so in particular if one of them is lost it will be replaced.
-Thus we can replace all four qubits in two rounds of error detection.
-Future work will investigate this paradigm by combining mid-circuit measurement, qubit reuse and reinitialization, along with periodic atom cooling to remove a fundamental obstacle to neutral atom quantum computing by limiting the depth of circuit any atom must undergo.
+c, The last circuit fixes this, halfway, by switching the direction of the final two CNOT gates MBG (23); CBO+ (24).
+The last two data qubits are replaced with fresh qubits, so in particular if one of them is lost it will be replaced.
+Thus all four qubits can be replaced in two rounds of error detection.
 We design and execute experiments that repeatedly apply logical CZ gates on a $\llbracket 4,2,2\rrbracket$ code block.
@@
 Error detection with qubit refresh is inserted between each pair of CZ gates, as illustrated in Fig.
-13.
+14.
 Note that with $j+1$ logical CZ gates, there are $j$ rounds of error detection, with $5+8j$ CZ gates applied to $5+2j$ qubits.
-Figure 13: (a) Encoded circuit that applies two rounds of encoded CZ gates to ${|\overline{++}\rangle}$, encoded in the $\llbracket 4,2,2\rrbracket$ code, with two rounds of error detection.
-Together, the two error detection rounds refresh all four data qubits.
+Figure 14: (a) Circuit that applies two rounds of logical CZ gates to ${|\overline{++}\rangle}$, encoded in the $\llbracket 4,2,2\rrbracket$ code, with two rounds of error detection that refresh all four qubits.
 (b) The corresponding logical circuit.
 Since the CZ gates cancel out, both measurements should give $+$.
-(c) The baseline circuit that we run for comparison.
-The $X$ echo pulses cancel out coherent dephasing errors from the CZ gates and movement.
+(c) The baseline circuit run for comparison.
+The $X$ gates cancel out coherent dephasing errors from the CZ gates and movement.
 Both measurements should give --.
-This baseline performs better than applying the two CZ gates sequentially without leaving the interaction zone.
-We also ran a very similar experiment, except applying logical dual CZ gates, $(H\otimes H)\,\text{CZ}\,(H\otimes H)$, using $S_{X}\otimes S_{X}^{\dagger}\otimes S_{X}^{\dagger}\otimes S_{X}$.
-The encoded experiments can be compared against corresponding unencoded baselines.
+This baseline performs better than applying the two CZ gates without leaving the interaction zone.
+The encoded experiment can be compared against a corresponding unencoded baseline.
 Given the system architecture, there are a few reasonable baselines from which to choose.
 For example, we may choose to return qubits to the storage zone between each CZ, or not.
-We have chosen to select unencoded baselines in which, as shown in Fig.
-13(a), between each pair of CZs or dual CZs, both qubits are moved back to the storage zone and Pauli $X$ or $Z$ gates, respectively, are applied to each of them to cancel coherent dephasing from the movement and two-qubit gates.
-These baselines performed the best of those that we tried, disadvantaging any comparison between encoded and unencoded.
-Results for both repeated logical CZ and repeated logical dual CZ, for up to 10 logical operations and 9 rounds of loss correction, are shown in Fig.
-14, with detailed data in Table 6.
-Since the two physical baselines compile to the same circuit, we have combined them into a single baseline for comparison.
+We have chosen to select a baseline in which, as shown in Fig.
+14(a), between each pair of CZs, both qubits are moved back to the storage zone and Pauli $X$ gates are applied to each of them to cancel coherent dephasing from the movement and two-qubit gates.
+This baseline performed the best of those that we tried, disadvantaging any comparison between encoded and unencoded.
+Results for up to 10 logical operations and 9 rounds of loss correction, are shown in Fig.
+15.
 The encoded repeated logical CZ experiments largely beat the corresponding baselines, including, for example, an encoded circuit consisting of 10 logical CZs with 9 loss-correction rounds outperforming an unencoded sequence of 10 CZs.
-Surprisingly, the repeated dual CZ experiment performs much worse for depths $8$ and $10$.
-Since it should be similar to repeated CZs, this could be from a mistake in scheduling the circuits, or possibly from machine parameter drift.
-Figure 14: Error rates for $\llbracket 4,2,2\rrbracket$ code repeated logical CZ and repeated logical dual CZ experiments, and the corresponding physical baseline, conditioned on no initial loss.
+Figure 15: Error rates for $\llbracket 4,2,2\rrbracket$ code repeated logical CZ experiments, and the corresponding physical baseline, conditioned on no initial loss.
 Error detection between logical gates (Fig.
-13) permits mid-circuit correction of atom loss.
-Table 6: Data for our $\llbracket 4,2,2\rrbracket$ repeated controlled-$Z$ and repeated dual controlled-$Z$ experiments.
-The encoded experiment with CZ or dual CZ depth $k+1$ has $k$ loss-correction rounds.
-It uses $5+2k$ physical qubits, with $5+8k$ physical CZ gates.
-A trial is prerejected if the cat state verification fails, and it is postrejected if an uncorrectable error is detected.
-Results for all available data, and for data in which there is no initial loss are reported separately.
-The unencoded baseline experiments, the same for CZ as for dual CZ, are on two qubits and include echos to cancel coherent errors.
-Trials with initial loss are prerejected, and those with circuit loss are postrejected.
-Restricting to trials with no initial loss has no effect on the baseline.
-      All data
-CZ or dual Loss-correction
-CZ depth rounds Trials Prerejected Postrejected $\Pr[\text{accept}]$ Errors Error rate
-Encoded CZs 2 1 3400 189 732 $73_{-2}^{+1}\%$ 34 $1.4^{+0.5}_{-0.4}\%$
-4 3 4600 587 2030 $43(1)\%$ 43 $2.2^{+0.7}_{-0.6}\%$
-6 5 3700 290 2408 $27(1)\%$ 41 $4(1)\%$
-8 7 3700 302 2674 $20(1)\%$ 37 $5^{+2}_{-1}\%$
-10 9 3700 318 2979 $10.9(1)\%$ 21 $5(2)\%$
-Encoded dual CZs 2 1 5120 633 1051 $67(1)\%$ 53 $1.5(0.4)\%$
-4 3 8640 1388 3340 $45(1)\%$ 104 $2.7(0.5)\%$
-6 5 5120 1156 2899 $21(1)\%$ 69 $6^{+2}_{-1}\%$
-8 7 5120 1270 3544 $7.7^{+0.8}_{-0.7}\%$ 63 $16^{+4}_{-3}\%$
-10 9 5120 1640 3242 $4.7(6)\%$ 38 $16^{+5}_{-4}\%$
-Baseline 2 4960 293 302 $88.0(9)\%$ 77 $1.8(4)\%$
-4 4960 200 431 $87.3(9)\%$ 113 $2.6_{-0.4}^{+0.5}\%$
-6 4960 202 566 $84.5(10)\%$ 186 $4.4_{-0.6}^{+0.7}\%$
-8 4660 202 654 $82(1)\%$ 210 $5.5_{-0.7}^{+0.8}\%$
-10 4960 284 860 $77(1)\%$ 374 $9.8_{-0.9}^{+1.0}\%$
-      Trials with no initial loss
-CZ or dual Loss-correction
-CZ depth rounds Trials Prerejected Postrejected $\Pr[\text{accept}]$ Errors Error rate
-Encoded CZs 2 1 3119 127 603 $77_{-2}^{+1}\%$ 21 $0.9^{+0.4}_{-0.3}\%$
-4 3 3787 197 1692 $50(2)\%$ 28 $1.5^{+0.6}_{-0.5}\%$
-6 5 3209 205 2048 $30(2)\%$ 32 $3(1)\%$
-8 7 2890 188 2026 $23(2)\%$ 29 $4^{+2}_{-1}\%$
-10 9 3004 210 2417 $13(1)\%$ 15 $4(2)\%$
-Encoded dual CZs 2 1 4282 333 746 $75(1)\%$ 31 $1.0^{+0.4}_{-0.3}\%$
-4 3 6732 854 2324 $53(1)\%$ 58 $1.6^{+0.5}_{-0.4}\%$
-6 5 4064 793 2302 $24(1)\%$ 48 $5(1)\%$
-8 7 3626 775 2500 $9.7^{+1.0}_{-0.9}\%$ 50 $14^{+4}_{-3}\%$
-10 9 3312 940 2171 $6.1(8)\%$ 26 $13^{+5}_{-4}\%$
-VI.2 Random circuits
+14) permits mid-circuit correction of atom loss.
+C.2 Random circuits
 As explained in Sec.
-III, the $\llbracket 4,2,2\rrbracket$ code allows a variety of other fault-tolerant logical gates.
-Each of $\{\mathrm{CNOT},\mathrm{CZ},\mathrm{SWAP},H^{\otimes 2}\}$ can be implemented by permutation or transversal gates.
-This set generates a group of size $36$.
+B, the $\llbracket 4,2,2\rrbracket$ code allows a variety of other fault-tolerant logical gates.
+Permutations and transversal gates can implement $\mathrm{CNOT},\mathrm{CZ},\mathrm{SWAP}$ and $H^{\otimes 2}$, generating a group of size $36$.
 To test the performance of these gates, and in the spirit of Ref.
@@
 Note that the error detection circuits used here are those of Fig.
-12(a), instead of those of Fig.
-12(b); therefore they correct for lost qubits only at the end of the circuit.
-Future work may include executing these logical operation sequences with the circuits of Fig.
-12(b).
-The choice to error-detect between group elements is justified by imagining a computation with many [[4,2,2]] blocks.
+13(a), instead of those of Fig.
+13(b); therefore they correct for lost qubits only at the end.
+The choice to error-detect between group elements is justified by imagining a computation with many $\llbracket 4,2,2\rrbracket$ blocks.
 In that scenario, it is natural to write a circuit as alternating layers of CNOTs between blocks, and group elements within each block.
@@
 The encoded circuits can be directly compared to corresponding physical baselines that execute the same gate sequence.
-Results are shown in Fig.
-15, with more details in Table 7.
-We find the encoded circuits consistently outperform the corresponding physical unencoded baselines.
+As shown in Fig.
+16, the encoded circuits outperform the corresponding physical baselines.
 Figure 16: Error rates for $\llbracket 4,2,2\rrbracket$ code random logical circuits, conditioned on no initial loss. Sequence labels are from the sequences given in Eq (1).
-Table 7: Data for our $\llbracket 4,2,2\rrbracket$ random circuit experiments.
-The sequences of operations are specified in Eq.
-(1).
-The encoding of sequence $k$ includes $k$ rounds of error detection, $5+2k$ physical qubits and $5+8k$ physical CZ gates.
-For sequences 0, 1, and 2, the two stabilizer generators of the output each have weight one and are measured simultaneously in each experiment.
-An error is reported if either of the two generators have incorrect eigenvalue.
-The stabilizer generators of sequence 3 are $\{XZ,ZX\}$.
-The encoded circuit measures both of these generators in the same experiment via transversal $Y$-basis measurements.
-The physical baselines require separate experiments for each generator.
-Results for the two generators are therefore reported separately.
-All data is reported.
-Restricting to trials with no initial loss has no effect on the baseline.
-    All data
-Sequence Trials Pre. Post. $\Pr[\text{accept}]$ Errors Error rate
-Encoded 0 1200 66 $42$ $91(2)\%$ 6 $0.6^{+0.6}_{-0.3}\%$
-1 2800 119 $467$ $79_{-2}^{+1}\%$ 5 $0.2^{+0.4}_{-0.1}\%$
-2 2500 132 $737$ $65(2)\%$ 9 $0.6^{+0.4}_{-0.3}\%$
-$XZ$ 3 7200 595 $3292$ $46(1)\%$ 15 $0.5^{+0.3}_{-0.2}\%$
-$ZX$ 3 7200 595 $3292$ $46(1)\%$ 24 $0.7(3)\%$
-Baseline 0 1600 52 $51$ $94(1)\%$ 20 $1.3^{+0.7}_{-0.5}\%$
-1 1600 40 $55$ $94(1)\%$ 13 $0.9^{+0.6}_{-0.4}\%$
-2 1600 66 $47$ $93(1)\%$ 23 $1.6^{+0.7}_{-0.5}\%$
-$XZ$ 3 1800 87 $108$ $89(1)\%$ 50 $3.1^{+0.9}_{-0.8}\%$
-$ZX$ 3 1800 108 $134$ $87(2)\%$ 40 $2.6^{+0.9}_{-0.7}\%$
-Trials with no initial loss
-Sequence Trials Pre. Post. $\Pr[\text{accept}]$ Errors Error rate
-Encoded 0 1104 45 $35$ $93_{-2}^{+1}\%$ 3 $0.3^{+0.5}_{-0.2}\%$
-1 2645 95 $398$ $81_{-2}^{+1}\%$ 3 $0.1^{+0.2}_{-0.1}\%$
-2 2289 105 $605$ $69(2)\%$ 9 $0.6^{+0.5}_{-0.3}\%$
-$XZ$ 3 6315 425 $2702$ $50(1)\%$ 15 $0.5^{+0.3}_{-0.2}\%$
-$ZX$ 3 6315 425 $2702$ $50(1)\%$ 24 $0.8(3)\%$
-VII Repeated error correction with the distance-three Bacon-Shor code
-Running deep logical circuits in a neutral atom quantum computer will ultimately require repeated correction of both errors within the qubit subspace and atom loss from erasure conversion.
-In order to demonstrate this, we choose the distance-three $\llbracket 9,1,3\rrbracket$ Bacon-Shor code.
-This code has previously been implemented on trapped-ion systems EDN+ (21); HBC (24).
-A code distance greater than two is needed to correct errors other than loss.
-To our knowledge, the present work is the first demonstration of repeated error correction in neutral atoms.
-We employ a scheme similar to that of Steane and Knill Ste (97); Kni (05); PSR+ (24), leveraging its ability to correct atom loss; see Fig.
-16.
-In this work, the ancilla qubits are held until readout at the end of the circuit, so each round of error correction requires the use of additional qubits.
-However, in principle, this scheme can be combined with mid-circuit measurement and reset where readout and loss detection, combined with mid-circuit reloading, will enable arbitrary-depth circuits.
-[…]
-After one round of error correction, we find a logical error rate of $4.9^{+1.0}_{-0.9}\%$ and after two rounds the error rate is $8^{+2}_{-1}\%$.
-Table 8 gives more details.
-While these results suggest the opportunity to further improve physical error rates, they also serve as a demonstration of the mechanism to correct for both errors and loss.
-This can be seen in the analysis of the successful trials with one round of error correction, where on average $0.20$ erasures and at least $0.14$ errors were corrected in the readout of the final data block, whereas $0.40$ erasures and $0.50$ errors were corrected in the previous round of $X$ and $Z$ error correction, a higher average number of corrections.
-In the successful trials with two rounds of error correction, on average $0.22$ erasures and at least $0.13$ errors were corrected in the readout of the final data block (consistent with the one-round data), while $0.95$ erasures and $0.88$ errors were corrected in the previous error-correction rounds.
-In order to compare encoded and physical results, we must choose a meaningful physical baseline.
-From the standpoint of the encoded data, the logical circuit is a quantum memory, applying the logical identity operation.
-For a meaningful comparison, we show results for repeated teleportation of physical qubit states, following the pattern of repeated teleportation of the logical data in Fig.
-16(a).
-Table 8: Data for $\llbracket 9,1,3\rrbracket$ Bacon-Shor code repeated error correction (EC) experiments.
-The data for the encoded circuits includes all trials, including those with initial loss, and there is no preselection.
-When decoding logical $X$ ($Z$), a trial is postrejected if there are losses in two or more rows (columns) of the $3\times 3$ code, or if one row is lost and the other two disagree.
-For the unencoded baseline, a trial is prerejected if there are any initial losses and a surviving trial is postrejected if there are any circuit losses.
-  Physical Physical
-Experiment qubits CZs Trials Prerejected Postrejected $\Pr[\text{accept}]$ Errors Error rate
-Encoded 1 EC round 27 36 2400 541 $77(2)\%$ 91 $4.9^{+1.0}_{-0.9}\%$
-2 EC rounds 45 66 2400 1054 $56(2)\%$ 114 $8^{+2}_{-1}\%$
-Baseline 2 1Q teleportations 3 2 $640$ 31 47 $88^{+2}_{-3}\%$ 7 $1.3^{+1.2}_{-0.7}\%$
-4 1Q teleportations 5 4 $640$ 50 82 $79(3)\%$ 7 $1.4^{+1.3}_{-0.8}\%$
-VIII Outlook
-In this work, we have created large entangled logical states, implemented computational algorithms on encoded qubits, demonstrated qubit loss correction, and implemented a code with the capacity for repeated loss and error correction.
-Altogether, these results highlight the unique features and promise of neutral atoms as a platform for reliable fault-tolerant quantum computing.
-The large encoded cat states and the Bernstein-Vazirani algorithms are enabled by the combination of scalability, high-fidelity operations inherent to neutral atoms, and all-to-all connectivity among the qubits.
-For larger-scale logical computation, this connectivity paired with high qubit count will enable the use of nonlocal error-correcting codes with high encoding rates and low physical qubit overhead XZZ+ (24).
-All of the encoded experiments, particularly those with repeated error detection, highlight the usefulness of loss-to-erasure conversion in the context of encoded qubits.
-They demonstrate that logical circuits can operate effectively with atom loss, which will be a key feature of large-scale neutral atom quantum computers.
-This work and similar work with other neutral atom platforms BEG+ (24), as well as recent work with superconducting qubits GC (24) and trapped ion qubits PSR+ (24); RAC+ (24), highlight a critical advance in the field of quantum computation: the transition from scientific results computed with physical qubits to computational advances achieved with logical qubits.
-After a relatively short time the neutral atom platform has shown significant advancement from its inception as an experimental quantum computing platform.
-In the near future we expect advances in both 2Q gate fidelities as more iterations are performed to reduce gate laser noise as well as significant scale up towards 10 0001000010\,00010 000 qubits allowing for larger distance codes, more performant logical qubits, and more impressive demonstrations of fault-tolerant, universal quantum computation.
-The combination of a hardware-optimized qubit virtualization system, complete with high-rate and high-efficiency quantum error correction, with a programmable neutral atom quantum processor, inclusive of mid-circuit measurement NC+ (23), continuous atom reloading NKC+ (24), and increased two-qubit gate fidelity, is a promising path to achieving large-depth, logical computation for scientific quantum advantage.
-Acknowledgments
-We thank Zulfi Alam, David Bohn, Jeongwan Haah, Chetan Nayak, Pasi Kostamo, Vadym Kliuchnikov, Dennis Tom, Matthias Troyer, and Jason Zander for their support and feedback during this project.
+Appendix D State preparation with the $\llbracket 16,6,4\rrbracket$ code
+Figure 17: a, Fault-tolerant preparation of encoded $|{+}\rangle^{\otimes 3}$ in the $\llbracket 8,3,2\rrbracket$ code.
+The blue part uses the gadget in Fig.
+13(b) to verify two of the stabilizers and swap qubits 3, 4 with 8, 9.
+The prepared state on qubits 0 to 7 is accepted only if qubits 8 and 9 are present and have trivial measurement outcomes.
+b, Non–fault-tolerant preparation.
+c, Layout of 250 atoms for preparing 25 copies of encoded $|{+}\rangle^{\otimes 3}$ using a.
+Colors only indicate the partition.
+Each $4\times 2$ tile joins a nearest same colored $2\times 1$ tile.
+d, Layout of 256 atoms for preparing 32 copies of encoded $|{+}\rangle^{\otimes 3}$ using b.
+e, Histogram of number of correct blocks of encoded $|{+}0{+}0{+}0\rangle$ in the $\llbracket 16,6,4\rrbracket$ code per shot, including shots with both $X$ and $Z$ basis measurements.
+Vertical lines indicate averages of 3.85(1) and 11.44(3) for the fault-tolerant and non–fault-tolerant protocol, respectively.
+We use all the 256 atoms to prepare error-corrected logical qubits encoded in blocks of the $\llbracket 16,6,4\rrbracket$ tesseract code RAC+ (24).
+To this end, we first prepare multiple copies of $|{+}\rangle^{\otimes 3}$ encoded in the $\llbracket 8,3,2\rrbracket$ color code Cam (16), and then randomly pair them up, each pair being a copy of $|{+}0{+}0{+}0\rangle$ encoded in the tesseract code RAC+ (24).
+We consider both fault-tolerant and non–fault-tolerant preparations of encoded $|{+}\rangle^{\otimes 3}$ in the $\llbracket 8,3,2\rrbracket$ code, as in Fig.
+17(a) and Fig.
+17(b), respectively.
+Compared to Fig.
+17(b), Fig.
+17(a) uses two extra qubits to measure the values of two stabilizers of the encoded $|{+}\rangle^{\otimes 3}$, and rejects upon any nontrivial measurement outcome or detected loss.
+For the non–fault-tolerant protocol, we first prepare 32 copies of encoded $|{+}\rangle^{\otimes 3}$, using the layout in Fig.
+17(d), and then randomly pair them up into 16 copies of encoded $|{+}0{+}0{+}0\rangle$.
+For the fault-tolerant protocol, we instead prepare 25 copies of encoded $|{+}\rangle^{\otimes 3}$ (with six atoms unused), as shown in Fig.
+17(c), and only pair up accepted ones.
+In both protocols, we measure all the atoms of the prepared tesseract code blocks in the same $X$ or $Z$ basis, and then decode.
+Specifically, an encoded $|{+}0{+}0{+}0\rangle$ block can be either “rejected” due to $\geq 2$ errors or losses detected, or “correct” if all the three deterministic values of logical $X$ or $Z$ operators are trivial, or “wrong” otherwise.
+For the fault-tolerant protocol, we run 1980 and 2000 trials for $X$ and $Z$ basis, respectively; for the non–fault-tolerant protocol, we run 960 and 1000 trials, respectively.
+For each of these trials, we repeat the random pairing of the (accepted) blocks of encoded $|+\rangle^{\otimes 3}$ independently for 100 shots.
+Results of these shots are shown in Fig.
+17(e).
+Specifically, we calculate the error rate, defined as the ratio of the number of wrong blocks over the number of correct or wrong blocks, both numbers summing over all shots of $X$ and $Z$ bases.
+We observe an error rate of $0.53(1)\%$ for the fault-tolerant protocol, nearly one order of magnitude smaller than $3.98(3)\%$ for the non–fault-tolerant protocol.
+However, the average number of correct blocks per shot is 3.85(1) for the fault-tolerant protocol, compared to 11.44(3) for the non–fault-tolerant protocol.
+This demonstrates a tradeoff between error rate and acceptance rate enabled by loss conversion.